### v1

# Exploring Model Kinship for Merging Large Language Models

Abstract
Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs). However, our understanding of the expected performance gains and principles when merging any two models remains limited. In this work, we introduce model kinship, the degree of similarity or relatedness between LLMs, analogous to biological evolution. With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models. Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can yield better performance on benchmark datasets. Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
1 Introduction
Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023). However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang & Yang, 2022; Lu et al., 2024; Liu et al., 2024). Recent studies suggest that model merging (Singh & Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena & Raffel, 2022; Yang et al., 2024a) offers a viable approach for achieving multitask objectives by integrating multiple expert models. Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
To date, through model merging techniques, resercheres have developed many more powerful LLMs through iterative model merging (Beeching et al., 2023), and to some extent, achieved model evolution (Figure 1(c)). Despite these successes, progress has predominantly relied on trial and error, along with extensive human expertise, but lacks formalized guidance and standardized procedures. As the merging iterations progress, achieving further generalization gains becomes increasingly challenging (More details in Section 3). For example, as shown in Figure 1, model merging often resembles the process of hybrid evolution in biology, where the next generation may not show significant improvements or may even regress, highlighting the imperative for a deeper exploration of the underlying mechanisms driving these advancements.
To address this, we introduce model kinship, a metric inspired by the concept of kinship (Sahlins, 2013) from evolutionary biology (Figure 1(a)). This metric is designed to estimate the degree of similarity or relatedness between LLMs during the iterative model merging process, offering insights intended to enhance the effectiveness of the merging strategy. We utilize the model kinship to conduct a comprehensive analysis of model merging experiments from two perspectives: the overall merging process, including various independent merge experiments and the evolution path of specific models, demostrating the complete merging trajectory.
Figure 1: An intuitive comparison between wheat evolution and model evolution. An interesting parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b). In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits. Similarly, in model merging, the merged model inherits parameters or weights from the contributing models. Part c demonstrates the iterative execution of model evolution. Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration. To be noted, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
Model kinship correlates with average performance gain in model merging. Emperically, we find that there is a strong correlation between variations in multitask capability, estimated by average task performance, and model kinship, which can help guide our selection of candidate models. We also notice that the model merging process has two stages: learning, where models experience significant performance improvements and saturation, where further improvements diminish and eventually stagnate. We think that the stagnation of improvements may be due to convergence in weight space, suggesting the presence of optimization challenges like local optima traps.
Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship. Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping to mitigate degradation and avoid local optima during model evolution. Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process. Overall, this paper makes three key contributions:
  1. 1.
    Introduction of Model Kinship: We introduce model kinship, designed to assess the degree of similarity or relatedness between LLMs during the merging process, which can guide model merging strategies and holds promise for advancing auto-merging research.
  2. 2.
    Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging. Our findings highlight the dynamics of multitask performance improvement and stagnation. Additionally, we propose a preliminary explaination of the underlying mechanisms using model kinship.
  3. 3.
    Practical Model Merging Strategies using Model Kinship: We demostrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
2 Background
2.1 Model Merging
Model merging aims to integrate two or more domain-specific models into a unified framework, thereby harnessing their compositive capabilities across multiple tasks (Sung et al., 2023). While this approach shares conceptual similarities with ensemble methods (Dietterich et al., 2002; Dong et al., 2020; Jiang et al., 2023b), model merging generates a single, generalized model, avoiding the increased inference time associated with ensembles. Let $f_{i}$ represent the $i$-th model for merging, each with its unique parameters $\theta_{i}$. If the merging process follows method $\mathcal{F}$, the prediction $\hat{y}$ of the merged model $f_{\text{merge}}$ for input $x$ is:
$\hat{y}=f_{\text{merge}}(x)=\mathcal{F}\left(f_{1}(x;\theta_{1}),f_{2}(x;\theta_{2}),\ldots,f_{n}(x;\theta_{n})\right)$ (1)
When employing parameter averaging methods, the merged model preserves the same architecture and parameter size as the original models, enabling it to be reused in future merging processes [^2] This paper focuses solely on model merging with identical architectures, where the merged model retains the same parameters as before the merge.. This makes it feasible to iteratively enhance the model through repeated applications of model merging. In this paper, we refer to this process as “Model Evolution”. Empirical evidence on open LLM leaderboard (Beeching et al., 2023) demonstrates that model evolution can successfully produce highly generalized models, often outperforming their original ones (Maxime Labonne, 2024).
2.2 Model Kinship
Intuitively, understanding the task capabilities of models is crucial for effective model merging. However, the black-box nature of LLMs presents significant challenges in comprehending their advantages Shi et al. (2024) and task-specific capabilities. For fine-tuned models, task benchmark results (Gao et al., 2024; Li et al., 2023c) and the characteristics of the fine-tuning datasets are often used as indicators of task-related knowledge. Yet, these metrics often fall short in comprehensive comparisons. The internal relationships between tasks and datasets can arise confusion, particularly in complex merging scenarios that involve multiple tasks.
Motivated by the parallel drawn between biological reproduction and the process of model evolution (as seen in Figure 1), we propose that the concept of kinship, which is central to evolutionary biology for understanding breeding relationships and human genealogies (Thompson, 1985), can be conceptually extended to the field of model merging, specifically to describe the kinship between models. To be noted, Ilharco et al. (2023) has proposed Task Arithmetic and offered new insights into how the merging process preserves and blends task information from the original models. (Ilharco et al., 2023) propose to use task vectors, defined as the differences in weight space between fine-tuned models and a pre-trained model, and demonstrate that arithmetic operations on these task vectors can serve as an effective method for model merging.
Inspired by this idea, we introduce the concept of Model Kinship, a metric designed to assess the degree of similarity or relatedness between large language models (LLMs) based on their “genetic” information (a.k.a. the changes in weights during model evolution). Considering two target models involved in a single merge, the weights of each target model are denoted as $\theta_{\text{target}}\in\mathbb{R}^{d}$. In this paper, it is significant to claim that a target model may either be a fine-tuned model or a previously merged model. Similarly, $\theta_{\text{base}}\in\mathbb{R}^{d}$ represents the weights of the base model from which the two target models are derived. The difference of weights $\delta_{\text{target}}$ between the target model and the base model is denoted as:
$\delta_{\text{target}}=\theta_{\text{target}}-\theta_{\text{base}}.$
For a model merging process where models have been fine-tuned or merged from the same base model, the model kinship $r$ can be defined using the similarity metric function $sim(\cdot,\cdot)$, which measures the similarity between two model weight differences. For $n$ models, we formally define the model kinship $r$ as:
$r=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),$
where $\delta_{\text{i}}$ and $\delta_{\text{j}}$ are the weight differences between the $i$-th and $j$-th target models and the base model, respectively.
3 Preliminary Analysis of Model Kinship
We begin with a preliminary analysis of model kinship with open-sourced LLMs from the community to better understand model evolution.
3.1 Evaluation Metrics
Let $T$ be the set of tasks in the task group, where $T=\{T_{1},T_{2},\ldots,T_{n}\}$. Each task $T_{i}$ in the set $T$ is associated with a performance measure $P_{i}$ for the LLM. For a multitask objective, the Average Task Performance (Avg.) $\bar{P}$ is calculated using the equation:
$\bar{P}=\frac{1}{n}\sum_{i=1}^{n}P_{\text{i}}$ (2)
To evaluate the effectiveness of a single merge, we propose the merge gain metric. Assume we have two models $m_{pre-1}$ and $m_{pre-2}$ and their average task performance are $\bar{P}_{pre-1}$ and $\bar{P}_{pre-2}$, intuitively, we believe the $\bar{P}_{\text{merged}}$ lie around the mean of $\bar{P}_{pre-1}$ and $\bar{P}_{pre-2}$. The merge gain is calculated as the difference of $\bar{P}_{\text{merged}}$ from the mean value of $\bar{P}_{pre-1}$ and $\bar{P}_{pre-2}$. For a merging recipe with $k$ models, the merge gain is:
$Gain=\bar{P}_{\text{merged}}-\frac{1}{k}\sum_{i=1}^{k}\bar{P}_{\text{pre-i}}$ (3)
In the following analysis, we use the task group T = {ARC, HellaSwag, MMLU, TruthfulQA, Winogrande, GSM8K}. All models are either fine-tuned or merged from the Mistral-7B architecture.
3.2 Correlation Analysis of Model Kinship and Performance Gain
Table 1: Correlation of Model Kinship based on different correlation function $sim(\cdot,\cdot)$ with Merge Gain, along with their corresponding p-values.
Metric Correlation Correlation
(Normal Value) (Absolute Value)
PCC -0.50 -0.59
P-value 0.063 0.023
CS -0.45 -0.66
P-value 0.098 0.008
ED 0.46 0.67
P-value 0.091 0.007
In this analysis, we examine the distribution of merge gain and model kinship based on Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED) in open-sourced LLMs, originating from the Mistral-7B (Jiang et al., 2023a). Those models are obtained from the HuggingFace, with assistance from the Open LLM Leaderboard (Details in Appendix A.).
3.2.1 Results
Figure 2: Distribution of Sample Experiments: Relationship Between Model Kinship (X-axis) and Merge Gain (Y-axis). Model Kinships are calculated using the Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED).
Figure 2 illustrates the distribution of model kinship based on three similarity metrics (PCC, CS, ED) in relation to merge gain. The scatter plots reveal a moderate correlation between model kinship and merge gain, as indicated by the trend lines. To further quantify these relationships, the correlation value (use Pearson Correlation Coefficient) between model kinship and merge gain are calculated, as detailed in the second column of Table 1. While moderate correlations are observed for all three metrics, the corresponding p-values indicate a weak level of statistical significance, ranging from 0.05 to 0.1. In contrast, when examining absolute merge gain, we find stronger and statistically significant correlations, as shown in the third column of Table 1. These results suggest that model kinship alone is insufficient for predicting whether a model can achieve generalization gains through merging. However, it may serve as a key factor in determining the upper limit of merge gains, thus indicating potential improvements. Since no significant differences are observed among the three metrics, we will focus solely on model kinship based on PCC in the following sections to simplify the demonstration.
3.3 Sequence Analysis of Model Evolution Paths
In this analysis, we examine changes in performance and model kinship across independent model evolution paths to identify the phased pattern of the merging process. We focus on the yamshadow experiment 28-7B (Labonne, 2024), a Mistral 7B architecture model ranked as the top 7B merged model on the Open LLM Leaderboard. From its model family tree, we extract two primary merging paths: Path 1 and Path 2.
3.3.1 Results
Figure 3: Change in Average Task Performance and Merge Gain across the Model Evolution process: The selected paths originate from two distinct initial models, with the saturation stage observed after the vertical line. Note that the generation of Path 2 is aligned with Path 1 for demonstration purposes.
Figure 4: Comparision between Model Kinship (measured by Pearson Correlation Coefficient) and Average Task Performance (normalized to the same value scale).
We first focus on the average task performance and merge gains throughout the model evolution path (Figure 3.) Detailed data and branch information are summarized in Appendix A). Our observations indicate that the performance improvements of the iterative merging process are not linear and can be divided into two stages:
  • Learning Stage. In this stage, the average task performance generally experiences a rapid increase. Noticeable merge gains suggest that the merged models are continually acquiring multitask capabilities through the merging process.
  • Saturation Stage. As the process continues, improvements begin to plateau. During this stage, the merge gains approach zero, indicating that the model can no longer benefit from the merging process and has ceased to improve.
Additionally, we compare the trend of model kinship with average task performance. Figure 4 illustrates the changes in model kinship alongside average task performance (normalized to the same range as the corresponding metric) throughout the model evolution paths. In addition to the notable correlation observed between average task performance and model kinship throughout the evolution process, model kinship exhibits a similar stage-specific pattern, particularly evident in the saturation stage, suggesting a potential relationship with the underlying cause of saturation.
3.4 Analysis of the Model Kinship in Different Merging Stages
Findings in previous analysis raises the question of “how model kinship affects performance improvements in model evolution”. To investigate the causality between model kinship and the stagnation of improvements, we examine the variation of model kinship across different merging stages from a broader perspective.
Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of top-performing models at each stage. For this analysis, we randomly select 5 models from each merging stage - Saturation Stage ($\geq 0.75$), Learning Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
Figure 5: The Model Kinship Matrices for the three model groups. Each element represents the model kinship value between the corresponding models. In Group B and C, the merged models are arranged by average task performance, ordered from high to low (left to right).
Figure 5 illustrates the model kinship between models within each group. Our analysis reveals that model kinship increases with the average task performance, even across models that follow different evolution paths yet share similar average performance levels. Additionally, during the saturation stage, all potential merges display a strong affinity, with model kinship values nearing 1. Since model kinship indicates the similarity of weights, we concluide the final findings as:
Findings: Model merging experiences a saturation stage, where the model kinship among top-performing models increases throughout the iterative merging process. This implies that the models converge to similar forms, resulting in excessive relatedness that undermines the effectiveness of the model merging strategy.
4 Using Model Kinship to Improve Model Merging
Inspired by the above findings, we further leverage model kinship to enhance the model merging process. We firstly conduct experiments employing a performance-prior greedy merging strategy. Note that the greedy strategy may eventually lead to convergence, we further introduce Top-$k$ Greedy Merging with Model Kinship (Algorithm 1). Our results indicate that while the greedy strategy focuses on short-term gains, it can lead to parameter convergence and suboptimal outcomes. By integrating model kinship, we can help the strategy avoid local optima. Furthermore, we find that model kinship holds potential for enhancing merging strategies as an early stopping criterion.
4.1 Experiment Setup
Algorithm 1 Top $k$ Greedy Merging with Model Kinship.
0:  A set $M$ of $n$ foundation models $\{M_{1},M_{2},\dots,M_{n}\}$, Evaluation function $f$, Similarity metric function $sim(\cdot,\cdot)$ for model kinship.
1:  Generate the first generation of merged models by merging each pair in set $M$$\{M_{1},M_{2},\dots,M_{n}\}$.
2:  Evaluate each merged model using $f$ and select the top $k$ models. Denote this set as $S=\{M_{1},M_{2},\dots,M_{n}\}$.
3:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $m$ models from the previous iteration.
4:  while $S\neq S_{\text{prev}}$ do
5:     Set $M_{\text{prev}}=M$.
6:     Set $S_{\text{prev}}=S$.
7:     Select $k$ pairs of models from $S$ with the highest performance according to $f$.
8:     Identify the current best model $M_{best}\in S$.
9:     Identify the model $M_{f}\in S$ with the highest model kinship to $M_{best}$ from the $M_{\text{prev}}$ according to the similarity metric $sim(\cdot,\cdot)$.
10:     Merge $M_{f}$ with $M_{best}$ to generate a new model $M_{\text{exp}}$ and add $M_{\text{exp}}$ into set $M$.
11:     Merge each selected pair to $M_{merged}$ (named as Model-gen-id) for merged models and add merged models into set $M$.
12:     Evaluate each new model using $f$ and update $S$ to be the new top $k$ models.
13:  end while
Note: The blue-highlighted steps are only executed in modified experiments incorporating model kinship-based exploration.
Figure 6: Left (a): The comparison of task performance improvement across merging generations. The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities. Right (b): The partial model family tree from the controled experiments. The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
LLMs.
We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210. Each model demonstrates distinct strengths across three benchmarks: Winogrande (5-shot), GSM8k (5-shot), and TruthfulQA MC2 (0-shot). All models are evaluated using the Language Model Evaluation Harness (Gao et al., 2024), a widely adopted framework for testing LLMs.
Merging Method.
We conduct two iterative model merging experiments, both utilizing the SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for merging [^3] We provide a Google Colab Notebook.. For implementation, we employ Mergekit (Goddard et al., 2024), a comprehensive toolkit that offers simple access to state-of-the-art model merging techniques.
Top $k$ Greedy Merging.
This strategy utilizes the vanilla Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1). This approach has been widely adopted in the community and has demonstrated notable success. In Figure 6 (b), models generated by the greedy strategy are indicated in green, while the best-performing models are highlighted in red.
Top $k$ Greedy Merging with Model Kinship.
The propposed strategy simply introduces an additional exploration step, based on model kinship, to the original greedy strategy (highlighted by the blue part in Algorithm 1). This approach aims to merge the best-performing model with the model that has the most distinct task capabilities, in order to discover potentially better solutions. In Figure 6 (b), models generated by our strategy are marked in purple, while the best-performing models are marked in yellow.
4.2 Results and Discussion
Figure 6 (a) illustrates the improvements in top average task performance across merging generations. Table 2 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
Both strategies achieve the multitask goals. However, the vanilla greedy strategy stops improving after Generation 2, stabilizing at an average task performance of 68.72. In contrast, Experiment 2, utilizing model kinship-based exploration, escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
Table 2: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach. The first three models serve as the foundation models in both experiments. Merged models are labeled using the structure: Model-{generation number}-{model identification number}.
Greedy Strategy + Model Kinship
Model Avg. Gain Kinship Model Avg. Gain Kinship
MetaMath 63.72 / / MetaMath 63.72 / /
Instruct 61.82 / / Instruct 61.82 / /
Open-chat 66.28 / / Open-chat 66.28 / /
model-1-1 62.17 -0.6 0.01 model-1-1 62.17 -0.6 0.01
model-1-2 64.02 -0.03 -0.02 model-1-2 64.02 -0.03 -0.02
model-1-3 66.84 +1.84 0.05 model-1-3 66.84 +1.84 0.05
model-2-1 68.72 +2.16 0.93 model-2-1 68.72 +2.16 0.93
model-2-2 61.47 -3.96 0.57 model-2-2 61.47 -3.96 0.57
model-2-3 61.32 -3.83 0.58 model-2-3 61.32 -3.83 0.58
model-3-1 68.59 +1.09 0.95 model-3-2 67.74 +1.09 0.93
model-3-2 67.74 -0.04 0.93 model-3-3 69.06 +0.74 0.24
- - - model-3-4 68.61 +1.13 0.32
model-4-1 68.51 -0.14 0.98 model-4-4 68.75 -0.14 0.54
model-4-2 68.04 -0.19 0.98 model-4-5 68.39 -0.27 0.66
model-4-3 68.53 +0.37 0.94 model-4-6 69.03 +0.15 0.52
- - - model-5-1 69.13 +0.04 0.65
- - - model-5-2 68.98 +0.07 0.65
- - - model-5-3 68.63 -0.37 0.98
Figure 7: Weight Change.
Merging Models with Low Kinship can Boost Exploration.
Figure 6(b) highlights the key branch of the model family tree. To investigate how merging models with low kinship helps escape local optima, we focus on the bifurcation point and analyze the weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3) in two separate experiments. The previous weight change, $v_{pre}$ (from Model-1-3 to Model-2-1), serves as a baseline. Figure 7 reveals that merging with the exploration model resulted in significant weight changes in a distinct direction, introducing novel variations into the weight space. In contrast, $v_{1}$ shows minimal weight change, as the merging effect is reduced due to the high similarity between the weights of openchat-3.5 and Model-2-1.
Early Stopping at High Kinship can Improve Efficiency.
We observe that the saturation stage of model evolution is particularly resource-intensive. In community experiments, 5 out of 14 merges in evolution path 1 resulted in an average improvement of just 0.57, while 3 out of 12 merges in evolution Path 2 yields an average improvement of 0.36. In our own experiments, applying a greedy strategy to a simple task lead to saturation after 2 out of 4 merges, with no further gains. These results indicate that human judgment and conventional stopping conditions cannot effectively halt the merging process at the optimal time. Therefore, we propose that model kinship can be used as an effective early stopping signal. When merging converges, the model kinship between top-performing models often exceeds 0.9. By halting the merging process at this point, time efficiency improves by approximately 30%, with minimal or no reduction in performance.
5 Related work
Weight averaging is one of the most widely used techniques in model merging, with its origins traced back to Utans (1996), who first applied it in neural networks to achieve performance comparable to ensemble methods. Since the 2010s, weight averaging has found numerous applications in deep neural networks, including combining checkpoints to enhance the training process (Nagarajan & Kolter, 2019; Tarvainen & Valpola, 2017; Izmailov et al., 2018; Li et al., 2023b; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023a; Smith & Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022). Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022). Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena & Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024). Ilharco et al. (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base. They demonstrate that arithmetic operations on these vectors enable model editing, such as achieving multitask learning. Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights. The Model Breadcrumbs (Davari & Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging. Merging models with different initializations requires additional considerations. Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh & Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024). While weight averaging cannot be directly applied to models with different architectures, it can still be used to enhance feasible fusion methods. Recent work, such as FuseChat (Wan et al., 2024b), combines weight averaging with Knowledge Fusion (Wan et al., 2024a) to develop innovative fusion techniques.
Recently, there have been some works exploring “model evolution”. Tellamekala et al. (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models. Labonne (2024) develop a tool to automatically merge models on Hugging Face, using an “Automerge” experiment to explore metrics in the merging process. Akiba et al. (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations, arguing that human intuition alone cannot uncover hidden patterns in merging.
6 Conclusion and Limitations
In this paper, we introduce model kinship, the degree of similarity or relatedness between LLMs, for merging LLMs, which can help guide our selection of candidate models. We conduct comprehensive experiments to demonstrate its effectiveness in understanding the model evolution process. We further propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship. We show that model kinship plays a crucial role in model evolution by guiding the process to escape local optima traps (in saturation stage), enabling further improvements. Additionally, we demonstrate that model kinship can detect the onset of convergence, allowing for early stopping and reducing the waste of computational resources in the merging process.
In a broad sense, our work explores how models can achieve autonomous evolution through model merging. Model merging can, to some extent, be likened to biological hybridization. Biological organisms have undergone billions of years of evolution to reach their current state. However, how silicon-based intelligence, represented by LLMs, evolves remains an unresolved mystery. We aspire that this work offer guidance and insights for the future merging and evolution of LLMs.
Limitations
However, there are several limitations to consider: a) The experiments in this study are conducted exclusively on models with the Mistral architecture, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu & Dao, 2023). b) The analysis relies on open-source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias. c) Correlation metrics for model kinship have not been fully explored. Other metrics may perform better than those discussed in this paper. d) The effectiveness of model kinship is demonstrated through empirical evidence. However, a theoretical framework (such as the assumptions in Appendix C) is needed to explain model evolution and model kinship more rigorously. e) Model kinship currently guides merging and improves performance limits but does not support sustained evolution. Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
Reproducibility Statement
The experimental setup can be found in Section 4.1. All model checkpoints are available on HuggingFace, with detailed information provided in Appendices A and B.
Appendix A Details of Experiments
All merged models from these experiments are accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets. The following tables cover two primary aspects:
  • (1) Information on the selected model family trees for two distinct evolution paths, along with detailed analysis results for each merge.
  • (2) A summary of the merge experiments conducted for distribution analysis.
The evolution paths are selected using a structured process, focusing on identifying key sequences within the model family trees. The steps were as follows:
  • Model Family Tree Construction: The complete model family tree is constructed by referencing model card details for each model involved.
  • Branch Identification: We identified the two longest branches within each tree, representing significant sequences of model merging.
  • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multitask capabilities and model kinship metrics.
Table 3 and 4 present detailed information on the sequential merging process. The second and third columns record the foundational models involved in each merge, while the final column lists the resulting merged models.
Table 3: Model Family tree of evolution Path 1.
Gen Model-1 Model-2 Model-Merged
1 Marcoroni-7B-v3 Mistral-7B-Merge-14-v0.1 distilabeled-Marcoro14-7B-slerp
2 distilabeled-Marcoro14-7B UNA-TheBeagle-7b-v1 Beagle14-7B
3 NeuralBeagle14-7B Turdus TurdusBeagle-7B
4 TurdusBeagle-7B FernandoGPT-v1 StrangeMerges_9-7B-dare_ties
5 StrangeMerges_9-7B-dare_ties MBX-7B-v3 StrangeMerges_10-7B-slerp
6 StrangeMerges_10-7B-slerp NeuralBeagle14-7B StrangeMerges_11-7B-slerp
7 StrangeMerges_11-7B-slerp MBX-7B-v3 StrangeMerges_20-7B-slerp
8 StrangeMerges_20-7B-slerp NeuTrixOmniBe-7B-model StrangeMerges_21-7B-slerp
9 StrangeMerges_21-7B-slerp Experiment26 StrangeMerges_30-7B-slerp
10 StrangeMerges_30-7B-slerp Experiment24 StrangeMerges_31-7B-slerp
11 StrangeMerges_31-7B-slerp Experiment28 StrangeMerges_32-7B-slerp
12 StrangeMerges_32-7B-slerp shadow-clown-7B-slerp
13 shadow-clown-7B-slerp yam-jom-7B YamShadow-7B
14 YamShadow-7B Experiment28 YamshadowExperiment28-7B
Table 4: Model Family tree of evolution Path 2.
Gen Model-1 Model-2 Model-Merged
1 neural-chat-7b-v3-3 openchat-3.5-1210 CatPPT-base
2 Marcoroni-7B-v3 CatPPT-base CatMacaroni-Slerp
3 LeoScorpius-7B CatMacaroni-Slerp SamirGPT-v1
4 SamirGPT-v1 Daredevil-7B
5 NeuralBeagle14-7B NeuralDaredevil-7B DareBeagle-7B
6 Turdus DareBeagle-7B TurdusDareBeagle-7B
7 MarcMistral-7B TurdusDareBeagle-7B MarcDareBeagle-7B
8 MarcBeagle-7B MarcDareBeagle-7B MBX-7B
9 MBX-7B pastiche-crown-clown-7b-dare
10 pastiche-crown-clown-7b-dare shadow-clown-7B-slerp
11 yam-jom-7B shadow-clown-7B-slerp YamShadow-7B
12 Experiment28-7B YamShadow-7B YamshadowExperiment28-7B
Table 5 and Table 6 present detailed analysis results that are not reported in the main paper. These include Average Task Performance (ATP), merge gains, and model kinship values, which are computed using Pearson Correlation coefficient, Cosine Similarity, and Euclidean Distance for each merge.
Table 5: Summary of Path 1 Results.
Gen Model-Merged ATP Gain PCC CS ED
1 distilabeled-Marcoro14-7B-slerp 73.63 0.55 0.82 0.76 5.15
2 Beagle14-7B 74.74 1.01 0.81 0.79 6.43
3 StrangeMerges_9-7B-dare_ties 75.15 0.45 0.93 0.89 4.66
4 StrangeMerges_9-7B-dare_ties 73.32 -0.69 0.90 0.84 4.70
5 StrangeMerges_10-7B-slerp 74.77 0.59 0.59 0.59 9.43
6 StrangeMerges_11-7B-slerp 74.8 0.045 0.87 0.86 5.31
7 StrangeMerges_20-7B-slerp 75.52 0.6 0.84 0.85 4.82
8 StrangeMerges_21-7B-slerp 76.29 0.38 0.85 0.89 4.28
9 StrangeMerges_30-7B-slerp 76.58 0.065 0.96 0.94 2.83
10 StrangeMerges_31-7B-slerp 76.67 -0.02 0.97 0.97 2.21
11 StrangeMerges_32-7B-slerp 76.68 0.11 0.99 0.99 0.62
12 shadow-clown-7B-slerp 76.64 -0.02 0.93 0.94 2.49
13 YamShadow-7B 76.6 -0.02 0.97 0.97 2.19
14 YamshadowExperiment28-7B 76.86 0.25 0.98 0.98 1.61
Table 6: Summary of Path 2 Results.
Gen Model-Merged ATP Gain PCC CS ED
1 CatPPT-base 72.25 2.89 0.02 0.01 20.41
2 CatMacaroni-Slerp 72.74 0.35 0.88 0.83 6.16
3 SamirGPT-v1 73.11 0.64 0.79 0.70 6.47
4 Daredevil-7B 74.12 0.33 0.87 0.83 4.81
5 DareBeagle-7B 74.58 0.15 0.79 0.77 6.01
6 TurdusDareBeagle-7B 74.94 0.32 0.90 0.86 4.59
7 MarcDareBeagle-7B 74.75 0.67 0.87 0.87 4.17
8 MBX-7B 75.04 0.11 0.96 0.96 2.90
9 pastiche-crown-clown-7b-dare 76.50 0.29 0.83 0.84 5.38
10 shadow-clown-7B-slerp 76.64 -0.02 0.93 0.94 2.49
11 YamShadow-7B 76.60 -0.02 0.97 0.97 2.19
12 YamshadowExperiment28-7B 76.86 0.25 0.98 0.98 1.61
Table 7 presents all merge experiments contributing to the distribution analysis. The selection of sample experiments adheres to two rules: (1) Samples are evenly chosen across average task performance values ranging from 0.7 to 0.7686 (the average task performance of the best 7B merged model) to accurately reflect the full scope of model evolution. (2) The experiments involve merges of two foundation models, as including multiple models introduces excessive noise.
Table 7: All Sample Experiments used in distribution analysis.
Model 1 Model 2 Merge Gain
Multi_verse_model-7B Experiment26-7B 0.06
M7-7b StrangeMerges_32-7B-slerp -0.03
Ognoexperiment27 Multi_verse_model-7B 0.03
YamShadow-7B Experiment28 0.25
shadow-clown-7B-slerp yam-jom-7B -0.02
StrangeMerges_21-7B-slerp Experiment26 0.06
StrangeMerges_31-7B-slerp Experiment28 0.11
NeuralBeagle14-7B Turdus 0.45
DareBeagle-7B Turdus 0.32
TurdusBeagle-7B FernandoGPT-v1 -0.69
StrangeMerges_10-7B-slerp NeuralBeagle14-7B 0.04
TurdusDareBeagle-7B MarcMistral-7B 0.67
StrangeMerges_20-7B-slerp NeuTrixOmniBe-7B-model-remix 0.38
StrangeMerges_11-7B-slerp MBX-7B-v3 0.6
Marcoroni-7B-v3 Mistral-7B-Merge-14-v0.1 0.55
distilabeled-Marcoro14-7B-slerp UNA-TheBeagle-7b-v1 1.01
UNA-TheBeagle-7b-v1 distilabeled-Marcoro14-7B-slerp 0.89
CatPPT-base Marcoroni-7B-v3 0.35
CatMacaroni-Slerp LeoScorpius-7B 0.64
NeuralDaredevil-7B NeuralBeagle14-7B 0.15
StrangeMerges_9-7B-dare_ties MBX-7B-v3 0.59
mistral-ft-optimized-1218 NeuralHerems-Mistral-2.5-7B -0.85
neural-chat-7b-v3-2 OpenHermes-2.5-Mistral-7B 1.91
StrangeMerges_30-7B-slerp Experiment24 -0.02
openchat-3.5-1210 neural-chat-7b-v3-3 2.89
MultiverseEx26-7B-slerp CalmExperiment-7B-slerp -0.09
CapybaraMarcoroni-7B DistilHermes-2.5-Mistral-7B 0.47
Multi_verse_model-7B Calme-7B-Instruct-v0.9 0.04
StrangeMerges_16-7B-slerp coven_7b_128k_orpo_alpha -0.35
Kunoichi-DPO-v2-7B AlphaMonarch-7B -1.05
StrangeMerges_15-7B-slerp Kunoichi-7B 0.39
Mistral-T5-7B-v1 Marcoroni-neural-chat-7B-v2 -0.18
Marcoro14-7B-slerp mistral-ft-optimized-1218 -0.61
mistral-ft-optimized-1218 NeuralHermes-2.5-Mistral-7B -0.85
MarcDareBeagle-7B MarcBeagle-7B -0.07
MetaMath-Mistral-7B Tulpar-7b-v2 -0.29
YugoGPT AlphaMonarch-7B -5.96
Appendix B Details of Model Group Selection
This appendix presents the exact models included in each model group, as shown in Table 8. The selection process is conducted across three distinct groups: (1) the top 5 models on the leaderboard, with a performance difference of 0.2, (2) 5 models with performance scores around 73 and a performance difference of 0.2, and (3) 5 fine-tuned models. It is important to note that the fine-tuned models were not selected based on performance, and may exhibit significant differences in results.
Table 8: Model Group in Kinship Matrix Analysis.
Group Models
Top Model Group YamshadowExperiment28-7B
Yamshadow-7B
Experiment25-7B
StrangeMerges_24-7B-slerp
MonaTrix-v6
Mid Stage Model Group Daredevil-7B
CatMarcoro14-7B
Mayo
Calmesmol-7B-slerp
StrangeMerges_4-7B-slerp
Fine-tuned Model Group Zephyr-beta
MetaMath-Mistral-7B
Mistral-7B-Instruct-v0.2
openchat-3.5-1210
WizardLM-2
Appendix C Assumption of Continual Model Merging
Figure 8: An intuitive illustration of the optimization process assumption in model evolution, where models progressively converge towards the optimal model.
Our findings in the main paper offer a new perspective on model evolution through multiple merging. If the merging process can be improved using a common optimization strategy, it raises the question of whether the underlying mechanism mirrors this optimization problem. Thus, we hypothesize the following:
Hypothesis: The evolution process may be simplified to a binary search process for most weight-averaging-based model merging methods.
Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy. It shows that Merged Model 2 may be overlooked, despite its potential for better multitask performance.
Figure 8 illustrates the ideal scenario in our assumption where multiple merges produce a highly generalized model. For the generalization task $t$, the y-axis represents the model performance for task $t$ and the x-axis represents the model’s weight space. In early merging stages, models fine-tuned with different tasks exhibit significant weight space dissimilarity. The merging process averages these weight spaces, and the experiment conductor selects the better-merged models while discarding the inferior ones. In stage 2, the search area narrows and the improvements become stable, eventually leading to an optimized state in stage 3 when “saturation stage” occurs.
In this context, Model Kinship serves as a metric to quantify the weight space distance between two models, with a higher model kinship indicating a lower weight space distance. Following this assumption, our findings of the optimization problem in model evolution can be elucidated in Figure 9.
However, we currently lack sufficient evidence to validate this hypothesis. Future work is needed to explore this further.
Appendix D Other Information and Results
This section presents additional results that are not reported in the main paper but may be useful for those interested. Figure 10 compares average task performance and model kinship using alternative metrics. Table 9 provides information on the exploration models used in our experiments. Figure 11 depicts a more “realistic” representation of combining two features/capabilities in evolution, highlighting the challenges in the selection process.
Figure 10: Illustration of comparison between the correlation of Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED) with average task performance (Normalized to the same value scale).
Table 9: Results of Exploration Models in our modified strategy.
Model Avg. Gain Kinship
model-2-exp 68.11 +3.78 -0.02
model-3-exp 68.15 +2.70 0.03
model-4-exp 68.49 +3.42 0.48
model-5-exp 68.36 -0.09 0.53
Figure 11: An intuitive comparison between selective breeding and continual model merging. The left process demonstrates breeding a tall and frutful plant by selecting parents with the desired traits in an biological scenario. The right process shows developing a model with capabilities of coding and math through model evolution.

### v2

@@
 Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs).
-However, our understanding of the expected performance gains and principles when merging any two models remains limited.
-In this work, we introduce model kinship, the degree of similarity or relatedness between LLMs, analogous to biological evolution.
+The open-source community has driven model evolution by iteratively merging existing models.
+However, a principled understanding of the expected gains and underlying factors in model merging remains lacking.
+In this work, we examine model evolution through continual merging, analogous to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs.
 With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models.
@@
 Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
+Exploring Model Kinship for Merging Large Language Models
+Yedi Hu, Yunzhi Yao, Shumin Deng , Huajun Chen, Ningyu Zhangthanks: Corresponding Author. Zhejiang University   National University of Singapore, NUS-NCS Joint Lab, Singapore {zhangningyu}@zju.edu.cn
 1 Introduction
 Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023).
-However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang & Yang, 2022; Lu et al., 2024; Liu et al., 2024).
-Recent studies suggest that model merging (Singh & Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena & Raffel, 2022; Yang et al., 2024a) offers a viable approach for achieving multitask objectives by integrating multiple expert models.
-Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
-To date, through model merging techniques, resercheres have developed many more powerful LLMs through iterative model merging (Beeching et al., 2023), and to some extent, achieved model evolution (Figure 1(c)).
-Despite these successes, progress has predominantly relied on trial and error, along with extensive human expertise, but lacks formalized guidance and standardized procedures.
-As the merging iterations progress, achieving further generalization gains becomes increasingly challenging (More details in Section 3).
-For example, as shown in Figure 1, model merging often resembles the process of hybrid evolution in biology, where the next generation may not show significant improvements or may even regress, highlighting the imperative for a deeper exploration of the underlying mechanisms driving these advancements.
-To address this, we introduce model kinship, a metric inspired by the concept of kinship (Sahlins, 2013) from evolutionary biology (Figure 1(a)).
-This metric is designed to estimate the degree of similarity or relatedness between LLMs during the iterative model merging process, offering insights intended to enhance the effectiveness of the merging strategy.
-We utilize the model kinship to conduct a comprehensive analysis of model merging experiments from two perspectives: the overall merging process, including various independent merge experiments and the evolution path of specific models, demostrating the complete merging trajectory.
+However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang and Yang, 2022; Lu et al., 2024; Liu et al., 2024).
 Figure 1: An intuitive comparison between wheat evolution and model evolution.
@@
 Starting with a group of LLMs, the repository evolves through a Selection-Merge-Recycle iteration.
-To be noted, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
-Model kinship correlates with average performance gain in model merging.
-Emperically, we find that there is a strong correlation between variations in multitask capability, estimated by average task performance, and model kinship, which can help guide our selection of candidate models.
-We also notice that the model merging process has two stages: learning, where models experience significant performance improvements and saturation, where further improvements diminish and eventually stagnate.
-We think that the stagnation of improvements may be due to convergence in weight space, suggesting the presence of optimization challenges like local optima traps.
-Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship.
-Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping to mitigate degradation and avoid local optima during model evolution.
+Notably, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
+Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multitask objectives by integrating multiple expert models.
+Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
+To date, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023).
+Many of these models are created through a biologically inspired evolutionary process that involves iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)).
+Despite these successes, the current approach faces critical limitations.
+Progress often relies on trial and error and extensive human expertise, with little formal guidance or standardized procedure.
+To address this problem, we propose a strategic model evolution framework (Figure 1(c)) that leverages explicit strategies to guide the direction of model evolution toward improved performance.
+We show that even a simple greedy strategy can outperform baseline merging approaches.
+However, in the late stages of both the community experiments and the greedy strategy evolution, achieving further gains in generalization becomes increasingly difficult.
+To explore possible solutions, we introduce ’model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process.
+Model kinship is designed to quantify the degree of similarity or relationship between models throughout the iterative merging process.
+By providing a principled way to measure these relationships, model kinship offers crucial insights for refining merging strategies.
+We conduct a comprehensive analysis of model merging experiments based on model kinship.
+We observe that the community model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau.
+Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of the potential effectiveness of a merge.
+These findings inspire two main insights: (1) high-kinship merges can lead to performance stagnation, akin to inbreeding; (2) low-kinship merges carry greater risk but may yield larger gains and facilitate escape from local optima.
+Inspired by this, we propose a new continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship.
+[^2] https://github.com/zjunlp/ModelKinship..
+Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping mitigate degradation and avoid local optima during model evolution.
 Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process.
-Overall, this paper makes three key contributions:
+In general, this paper makes four key contributions:
   1. 1.
-    Introduction of Model Kinship: We introduce model kinship, designed to assess the degree of similarity or relatedness between LLMs during the merging process, which can guide model merging strategies and holds promise for advancing auto-merging research.
+    Continual Model Merging as a Feasible Framework for Model Evolution: We propose continual model merging as a viable framework for evolving large language models.
+Through strategically guided merging across iterations, this approach yields consistent improvements in generalization and task performance.
   2. 2.
+    Introducing Model Kinship: We introduce model kinship, designed to assess the degree of similarity or relationship between LLMs during the merging process, which can guide model merging strategies and holds promise for advancing auto merging research.
+  3. 3.
     Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging.
-Our findings highlight the dynamics of multitask performance improvement and stagnation.
-Additionally, we propose a preliminary explaination of the underlying mechanisms using model kinship.
-  3. 3.
-    Practical Model Merging Strategies using Model Kinship: We demostrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
+Our findings highlight the dynamics of multitask performance improvement and stagnation during evolution.
+In addition, we propose a preliminary explanation of the underlying mechanisms using model kinship.
+  4. 4.
+    Practical Model Merging Strategies using Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
 2 Background
-2.1 Model Merging
-Model merging aims to integrate two or more domain-specific models into a unified framework, thereby harnessing their compositive capabilities across multiple tasks (Sung et al., 2023).
+2.1 Model Merging: Fundamentals
+Model merging aims to integrate two or more domain-specific models into a unified framework, thereby harnessing their composite capabilities across multiple tasks (Sung et al., 2023).
 While this approach shares conceptual similarities with ensemble methods (Dietterich et al., 2002; Dong et al., 2020; Jiang et al., 2023b), model merging generates a single, generalized model, avoiding the increased inference time associated with ensembles.
@@
 $\hat{y}=f_{\text{merge}}(x)=\mathcal{F}\left(f_{1}(x;\theta_{1}),f_{2}(x;\theta_{2}),\ldots,f_{n}(x;\theta_{n})\right)$ (1)
-When employing parameter averaging methods, the merged model preserves the same architecture and parameter size as the original models, enabling it to be reused in future merging processes [^2] This paper focuses solely on model merging with identical architectures, where the merged model retains the same parameters as before the merge..
-This makes it feasible to iteratively enhance the model through repeated applications of model merging.
-In this paper, we refer to this process as “Model Evolution”.
-Empirical evidence on open LLM leaderboard (Beeching et al., 2023) demonstrates that model evolution can successfully produce highly generalized models, often outperforming their original ones (Maxime Labonne, 2024).
+2.2 Continual Model Merging: Benefits and Challenges
+Parameter averaging methods enable the merged model to retain the same architecture and parameter size as the original models, allowing for reuse in future merging processes.
+Leveraging this property, the community has progressively enhanced models through repeated applications of model merging, a process we refer to as “Model Evolution”.
+Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that model evolution can yield highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
+However, current model evolution practices typically involve random merging by multiple contributors, leading to high computational costs and unstable behavior that limit its feasibility for industrial applications.
+3 Strategic Model Evolution
+In this section, we demonstrate how strategic model evolution via continual merging can achieve better generalization across tasks.
+3.1 Method
+We first define an iterative process for model evolution, where models are incrementally merged to produce a more generalized model.
+The direction of model evolution is guided by a principled strategy that governs the model selection.
+At each iteration $t$, a selection strategy $\mathcal{S}$ is applied to the current pool of models $\mathcal{P}_{t}=\{M_{t}^{(1)},M_{t}^{(2)},\ldots,M_{t}^{(k)}\}$, producing a subset of candidate models $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$.
+These selected models are then combined using a merge method $\mathcal{M}$, such as weighted averaging or linear interpolation in the weight space, resulting in the next-generation model $M_{t+1}$.
+This iterative process continues until a predefined end strategy $\mathcal{E}$ is matched.
+Formally, the process is defined as:
+$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{subject to}\quad\neg\mathcal{E}(M_{t+1})$ (2)
+Method TruthfulQA Winogrande GSM8k Avg.
+Ties 62.76 78.61 0.11 47.16
+Linear 56.37 78.08 68.54 67.66
+Continual Merge (ord1) 47.15 76.24 53.15 58.84
+Continual Merge (ord2) 61.01 79.56 63.76 68.11
+Continual Merge (ord3) 49.80 78.53 55.72 61.35
+Strategic Model Evolution 50.94 80.11 75.13 68.72
+Table 1: Performance Comparison across different merging methods.
+3.2 Setup
+Baseline Method.
+We adopt two commonly used approaches for merging multiple models as baselines.
+The first is multi-model merging, where all models are merged simultaneously.
+The second is sequential continual merging, where three models are merged in sequence using a pairwise merging strategy.
+For these baselines, we apply TIES (Yadav et al., 2023) and Linear Merge for the multi-model merging case and SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for the pairwise merging in the sequential strategy.
+Top $k$ Greedy Merging.
+Our method applies continual model merging with greedy strategy, the Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1 in Appendix C).
+Each single merging step in the continual merging experiments is performed using SLERP.
+Models and Datasets.
+We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
+Evaluation.
+Evaluation is performed using three task-specific benchmark datasets: Winogrande, GSM8k, and TruthfulQA.
+These benchmarks demonstrate the distinct strengths of the three selected fine-tuned models.
+Further details on the tasks are provided in Appendix A.3.
+3.3 Results
+As shown in Table 1, Strategic Model Evolution can yield better generalization when combining multiple tasks.
+However, compared to simple sequential merging, the improvements are limited due to the simplicity of the greedy strategy.
+This observation raises an important question regarding optimal merging strategies:
+Problem: How can we design more principled and effective strategies for model merging to better facilitate model evolution?
+4 Preliminary Analysis of Model Kinship
+To explore efficient strategies for model evolution, we propose a novel metric, model kinship, which captures task-related differences between models and may help predict merging outcomes.
+We then conduct a preliminary analysis of model kinship in community merging experiments on LLMs to evaluate its potential for informing and enhancing model evolution strategies.
 4.1 Model Kinship
-Intuitively, understanding the task capabilities of models is crucial for effective model merging.
-However, the black-box nature of LLMs presents significant challenges in comprehending their advantages Shi et al.
-(2024) and task-specific capabilities.
-For fine-tuned models, task benchmark results (Gao et al., 2024; Li et al., 2023c) and the characteristics of the fine-tuning datasets are often used as indicators of task-related knowledge.
-Yet, these metrics often fall short in comprehensive comparisons.
-The internal relationships between tasks and datasets can arise confusion, particularly in complex merging scenarios that involve multiple tasks.
-Motivated by the parallel drawn between biological reproduction and the process of model evolution (as seen in Figure 1), we propose that the concept of kinship, which is central to evolutionary biology for understanding breeding relationships and human genealogies (Thompson, 1985), can be conceptually extended to the field of model merging, specifically to describe the kinship between models.
-To be noted, Ilharco et al.
-(2023) has proposed Task Arithmetic and offered new insights into how the merging process preserves and blends task information from the original models.
-(Ilharco et al., 2023) propose to use task vectors, defined as the differences in weight space between fine-tuned models and a pre-trained model, and demonstrate that arithmetic operations on these task vectors can serve as an effective method for model merging.
-Inspired by this idea, we introduce the concept of Model Kinship, a metric designed to assess the degree of similarity or relatedness between large language models (LLMs) based on their “genetic” information (a.k.a.
-the changes in weights during model evolution).
-Considering two target models involved in a single merge, the weights of each target model are denoted as $\theta_{\text{target}}\in\mathbb{R}^{d}$.
-In this paper, it is significant to claim that a target model may either be a fine-tuned model or a previously merged model.
-Similarly, $\theta_{\text{base}}\in\mathbb{R}^{d}$ represents the weights of the base model from which the two target models are derived.
-The difference of weights $\delta_{\text{target}}$ between the target model and the base model is denoted as:
-$\delta_{\text{target}}=\theta_{\text{target}}-\theta_{\text{base}}.$
-For a model merging process where models have been fine-tuned or merged from the same base model, the model kinship $r$ can be defined using the similarity metric function $sim(\cdot,\cdot)$, which measures the similarity between two model weight differences.
-For $n$ models, we formally define the model kinship $r$ as:
-$r=\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),$
-where $\delta_{\text{i}}$ and $\delta_{\text{j}}$ are the weight differences between the $i$-th and $j$-th target models and the base model, respectively.
-3 Preliminary Analysis of Model Kinship
-We begin with a preliminary analysis of model kinship with open-sourced LLMs from the community to better understand model evolution.
+Drawing inspiration from the parallel between artificial selection and model evolution (as detailed in Appendix G), we hypothesize that a concept analogous to kinship, the genetic relatedness studied in evolutionary biology (Thompson, 1985), can also apply to model merging.
+Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between models.
+This analogy suggests that just as genetic kinship influences breeding outcomes, model kinship may similarly impact the effectiveness of merging strategies in enhancing generalization performance.
+We adopt the most intuitive representation, inspired by the cosine similarity analysis introduced in the Task Arithmetic paper (Ilharco et al., 2023).
+This metric is designed to evaluate the degree of similarity or relatedness between the task capabilities of large language models (LLMs) based solely on their "genetic" information, meaning the changes in their weights, during model evolution.
+Considering two models $m_{i}$, $m_{j}$ involved in a model evolution originated from the pre-trained model $m_{base}$, the weights of $m_{i}$, $m_{j}$ are denoted as $\theta_{\text{i}},\theta_{\text{j}}\in\mathbb{R}^{d}$.
+Similarly, $\theta_{\text{base}}\in\mathbb{R}^{d}$ represents the weights of the pre-trained model.
+Since the differences between models emerge after fine-tuning and merging, the variation of weights during model evolution is crucial.
+It is calculated as:
+$\delta_{\text{i}}=\theta_{\text{i}}-\theta_{\text{base}},\delta_{\text{j}}=\theta_{\text{j}}-\theta_{\text{base}}$ (3)
+Model kinship $r$ is designed to capture the similarity of task capabilities between models.
+In this paper, we explore multiple potential metrics for evaluating similarity.
+For the calculation, $sim(\cdot,\cdot)$ denotes the similarity metric function used.
+Considering two cases merging of 2 models and merging of $n$ models, we formally define model kinship $r$ as:
+$r=\begin{cases}sim(\delta_{1},\delta_{2}),&\text{(Merge 2)}\\ \frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),&\text{(Merge N)}\end{cases}$ (4)
+We investigate the relationship between task performance and model kinship (see Appendix E for the full analysis).
+The results reveal strong correlations, reinforcing the view that model kinship reflects task-related differences between models.
 4.2 Evaluation Metrics
@@
 In the following analysis, we use the task group T = {ARC, HellaSwag, MMLU, TruthfulQA, Winogrande, GSM8K}. All models are either fine-tuned or merged from the Mistral-7B architecture.
-3.2 Correlation Analysis of Model Kinship and Performance Gain
-Table 1: Correlation of Model Kinship based on different correlation function $sim(\cdot,\cdot)$ with Merge Gain, along with their corresponding p-values.
+Figure 2: Distribution of Sample Experiments: Relationship Between Model Kinship (X-axis) and Merge Gain (Y-axis). Model kinships are calculated using PCC, CS, and ED.
+Figure 3: Change in Average Task Performance and Merge Gain across the Model Evolution Process: Paths originate from two different base models.
+The vertical line marks the transition to the saturation stage.
+Path 2 is temporally aligned with Path 1 for clarity.
+4.3 Analysis of Model Kinship: Correlation and Evolution Dynamics
+In this section, we analyze model kinship from two perspectives: (1) its correlation with performance gain across a broad range of open-sourced LLM merges, and (2) its dynamic along specific model evolution paths.
+These analyses aim to clarify the relationship between model kinship and multitask capability improvements, as well as to identify phases of merge effectiveness.
+4.3.1 Correlation Between Model Kinship and Performance Gain
+We begin by exploring the potential relationship between merge gain and model kinship using three similarity metrics: Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED).
+The models used in this analysis are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
+As illustrated in Figure 2, the scatter plots derived from all three metrics suggest a moderate correlation between model kinship and merge gain.
+Table 2 reports Pearson correlation values for both signed and absolute merge gains.
+While the correlations for signed gains appear relatively weak (with p-values between 0.05 and 0.1), those for absolute merge gains are comparatively stronger and show greater statistical significance.
+These observations imply that model kinship may offer some indication of the potential magnitude of merge gains, though it appears less effective at predicting the direction of change.
+While we cannot assert a causal relationship, the association provides useful insight into how kinship might relate to merge outcomes.
+In light of the comparable performance across the three metrics, we use PCC-based kinship in the remainder of our analysis for consistency.
+4.3.2 Model Kinship in Evolution Paths
+As a further exploration, we examine model kinship across independent model evolution paths to investigate potential phase patterns in the merging process.
+This analysis centers on the yamshadow experiment 28-7B (Labonne, 2024), a Mistral-7B-based model that ranks among the top-performing merged models on the Open LLM Leaderboard.
+From its model family tree, we extract two main merging trajectories, referred to as Path 1 and Path 2, for comparison.
+Figure 4: Comparison between Model Kinship (PCC) and Average Task Performance (normalized to the same scale).
+Figure 5: Left (a): The comparison of task performance improvement across merging generations.
+The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities.
+Right (b): The partial model family tree from the controlled experiments.
+The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
+Figure 3 displays the average task performance and the merge gains along the two evolution paths. The merging process exhibits two distinct phases:
+  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multitask balance.
+  • Saturation Stage. Performance stabilizes, and additional merges result in minimal or no measurable improvement.
+Figure 4 shows how the model kinship and normalized average performance change over the course of the merging process.
+Both metrics display a similar two-stage trend: they increase during the Improving Stage and level off during the Saturation Stage.
+This parallel trajectory suggests a potential link between model kinship and performance gains, indicating that improvements in generalization may coincide with and possibly depend on increases in model kinship.
 Metric Correlation Correlation
@@
 P-value 0.091 0.007
-In this analysis, we examine the distribution of merge gain and model kinship based on Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED) in open-sourced LLMs, originating from the Mistral-7B (Jiang et al., 2023a).
-Those models are obtained from the HuggingFace, with assistance from the Open LLM Leaderboard (Details in Appendix A.).
-3.2.1 Results
-Figure 2: Distribution of Sample Experiments: Relationship Between Model Kinship (X-axis) and Merge Gain (Y-axis).
-Model Kinships are calculated using the Pearson Correlation Coefficient (PCC), Cosine Similarity (CS) and Euclidean Distance (ED).
-Figure 2 illustrates the distribution of model kinship based on three similarity metrics (PCC, CS, ED) in relation to merge gain.
-The scatter plots reveal a moderate correlation between model kinship and merge gain, as indicated by the trend lines.
-To further quantify these relationships, the correlation value (use Pearson Correlation Coefficient) between model kinship and merge gain are calculated, as detailed in the second column of Table 1.
-While moderate correlations are observed for all three metrics, the corresponding p-values indicate a weak level of statistical significance, ranging from 0.05 to 0.1.
-In contrast, when examining absolute merge gain, we find stronger and statistically significant correlations, as shown in the third column of Table 1.
-These results suggest that model kinship alone is insufficient for predicting whether a model can achieve generalization gains through merging.
-However, it may serve as a key factor in determining the upper limit of merge gains, thus indicating potential improvements.
-Since no significant differences are observed among the three metrics, we will focus solely on model kinship based on PCC in the following sections to simplify the demonstration.
-3.3 Sequence Analysis of Model Evolution Paths
-In this analysis, we examine changes in performance and model kinship across independent model evolution paths to identify the phased pattern of the merging process.
-We focus on the yamshadow experiment 28-7B (Labonne, 2024), a Mistral 7B architecture model ranked as the top 7B merged model on the Open LLM Leaderboard.
-From its model family tree, we extract two primary merging paths: Path 1 and Path 2.
-3.3.1 Results
-Figure 3: Change in Average Task Performance and Merge Gain across the Model Evolution process: The selected paths originate from two distinct initial models, with the saturation stage observed after the vertical line.
-Note that the generation of Path 2 is aligned with Path 1 for demonstration purposes.
-Figure 4: Comparision between Model Kinship (measured by Pearson Correlation Coefficient) and Average Task Performance (normalized to the same value scale).
-We first focus on the average task performance and merge gains throughout the model evolution path (Figure 3.) Detailed data and branch information are summarized in Appendix A).
-Our observations indicate that the performance improvements of the iterative merging process are not linear and can be divided into two stages:
-  • Learning Stage.
-In this stage, the average task performance generally experiences a rapid increase.
-Noticeable merge gains suggest that the merged models are continually acquiring multitask capabilities through the merging process.
-  • Saturation Stage.
-As the process continues, improvements begin to plateau.
-During this stage, the merge gains approach zero, indicating that the model can no longer benefit from the merging process and has ceased to improve.
-Additionally, we compare the trend of model kinship with average task performance.
-Figure 4 illustrates the changes in model kinship alongside average task performance (normalized to the same range as the corresponding metric) throughout the model evolution paths.
-In addition to the notable correlation observed between average task performance and model kinship throughout the evolution process, model kinship exhibits a similar stage-specific pattern, particularly evident in the saturation stage, suggesting a potential relationship with the underlying cause of saturation.
-3.4 Analysis of the Model Kinship in Different Merging Stages
-Findings in previous analysis raises the question of “how model kinship affects performance improvements in model evolution”.
-To investigate the causality between model kinship and the stagnation of improvements, we examine the variation of model kinship across different merging stages from a broader perspective.
-Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of top-performing models at each stage.
-For this analysis, we randomly select 5 models from each merging stage - Saturation Stage ($\geq 0.75$), Learning Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
-Figure 5: The Model Kinship Matrices for the three model groups.
-Each element represents the model kinship value between the corresponding models.
-In Group B and C, the merged models are arranged by average task performance, ordered from high to low (left to right).
-Figure 5 illustrates the model kinship between models within each group.
-Our analysis reveals that model kinship increases with the average task performance, even across models that follow different evolution paths yet share similar average performance levels.
-Additionally, during the saturation stage, all potential merges display a strong affinity, with model kinship values nearing 1.
-Since model kinship indicates the similarity of weights, we concluide the final findings as:
-Findings: Model merging experiences a saturation stage, where the model kinship among top-performing models increases throughout the iterative merging process.
-This implies that the models converge to similar forms, resulting in excessive relatedness that undermines the effectiveness of the model merging strategy.
-4 Using Model Kinship to Improve Model Merging
-Inspired by the above findings, we further leverage model kinship to enhance the model merging process.
-We firstly conduct experiments employing a performance-prior greedy merging strategy.
-Note that the greedy strategy may eventually lead to convergence, we further introduce Top-$k$ Greedy Merging with Model Kinship (Algorithm 1).
-Our results indicate that while the greedy strategy focuses on short-term gains, it can lead to parameter convergence and suboptimal outcomes.
-By integrating model kinship, we can help the strategy avoid local optima.
-Furthermore, we find that model kinship holds potential for enhancing merging strategies as an early stopping criterion.
-4.1 Experiment Setup
-Algorithm 1 Top $k$ Greedy Merging with Model Kinship.
-0:  A set $M$ of $n$ foundation models $\{M_{1},M_{2},\dots,M_{n}\}$, Evaluation function $f$, Similarity metric function $sim(\cdot,\cdot)$ for model kinship.
-1:  Generate the first generation of merged models by merging each pair in set $M$$\{M_{1},M_{2},\dots,M_{n}\}$.
-2:  Evaluate each merged model using $f$ and select the top $k$ models. Denote this set as $S=\{M_{1},M_{2},\dots,M_{n}\}$.
-3:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $m$ models from the previous iteration.
-4:  while $S\neq S_{\text{prev}}$ do
-5:     Set $M_{\text{prev}}=M$.
-6:     Set $S_{\text{prev}}=S$.
-7:     Select $k$ pairs of models from $S$ with the highest performance according to $f$.
-8:     Identify the current best model $M_{best}\in S$.
-9:     Identify the model $M_{f}\in S$ with the highest model kinship to $M_{best}$ from the $M_{\text{prev}}$ according to the similarity metric $sim(\cdot,\cdot)$.
-10:     Merge $M_{f}$ with $M_{best}$ to generate a new model $M_{\text{exp}}$ and add $M_{\text{exp}}$ into set $M$.
-11:     Merge each selected pair to $M_{merged}$ (named as Model-gen-id) for merged models and add merged models into set $M$.
-12:     Evaluate each new model using $f$ and update $S$ to be the new top $k$ models.
-13:  end while
-Note: The blue-highlighted steps are only executed in modified experiments incorporating model kinship-based exploration.
-Figure 6: Left (a): The comparison of task performance improvement across merging generations.
-The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities.
-Right (b): The partial model family tree from the controled experiments.
-The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
-LLMs.
-We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
-Each model demonstrates distinct strengths across three benchmarks: Winogrande (5-shot), GSM8k (5-shot), and TruthfulQA MC2 (0-shot).
-All models are evaluated using the Language Model Evaluation Harness (Gao et al., 2024), a widely adopted framework for testing LLMs.
-Merging Method.
-We conduct two iterative model merging experiments, both utilizing the SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for merging [^3] We provide a Google Colab Notebook..
-For implementation, we employ Mergekit (Goddard et al., 2024), a comprehensive toolkit that offers simple access to state-of-the-art model merging techniques.
-Top $k$ Greedy Merging.
-This strategy utilizes the vanilla Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1).
-This approach has been widely adopted in the community and has demonstrated notable success.
-In Figure 6 (b), models generated by the greedy strategy are indicated in green, while the best-performing models are highlighted in red.
-Top $k$ Greedy Merging with Model Kinship.
-The propposed strategy simply introduces an additional exploration step, based on model kinship, to the original greedy strategy (highlighted by the blue part in Algorithm 1).
-This approach aims to merge the best-performing model with the model that has the most distinct task capabilities, in order to discover potentially better solutions.
-In Figure 6 (b), models generated by our strategy are marked in purple, while the best-performing models are marked in yellow.
-4.2 Results and Discussion
-Figure 6 (a) illustrates the improvements in top average task performance across merging generations.
-Table 2 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
-Both strategies achieve the multitask goals.
-However, the vanilla greedy strategy stops improving after Generation 2, stabilizing at an average task performance of 68.72.
-In contrast, Experiment 2, utilizing model kinship-based exploration, escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
-Table 2: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach.
-The first three models serve as the foundation models in both experiments.
-Merged models are labeled using the structure: Model-{generation number}-{model identification number}.
+Table 2: Correlation of Model Kinship based on different correlation function $sim(\cdot,\cdot)$ with Merge Gain, along with their corresponding p-values.
+Additionally, to generalize the findings from the Evolution Paths, we analyze how model kinship evolves across different stages of the merging process, thereby broadening the scope from individual paths to the full evolution picture.
+(detailed in Appendix D.1).
+The results show that the best performing models exhibit high mutual kinship, which may lead to a stagnation stage in the merging process.
+4.4 Discussion
+Considering all results that we observed, this analysis provides two main inspirations for the application of model kinship:
+  • High kinship merges may lead to performance stagnation, similar to biological inbreeding.
+  • Low kinship merges involve risk, but can lead to greater gains, potentially enabling escape from local optima caused by greedy strategy.
 Greedy Strategy + Model Kinship
@@
 - - - model-5-3 68.63 -0.37 0.98
-Figure 7: Weight Change.
+Table 3: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach.
+The first three models serve as the foundation models in both experiments.
+Note: The model kinship experiment was terminated at generation 5, as it has already outperformed the greedy strategy by that point.
+5 Using Model Kinship to Improve Model Merging
+Building on the insights from section 4, we explore how model kinship can be leveraged to improve the model merging process.
+Our main experiment centers on the Mistral-7B model, with detailed results presented in the main text.
+To further evaluate the generalization of our approach, we also conduct two supplementary experiments: one based on Llama-2 (see Appendix B) and another on a distinct task set to test robustness across different evaluation settings.
+Our results indicate that while greedy strategy focuses on short-term gains, it can lead to suboptimal outcomes.
+By integrating model kinship, we can help the strategy avoid local optima and gain further improvements.
+5.1 Main Experiment Setup
+For the main experiments, we follow the same settings as described in section 3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA.
+We adopt the Top-$k$ Greedy Merging strategy as the baseline continual merging strategy.
+Top $k$ Greedy Merging with Model Kinship.
+We propose an enhanced merging strategy that augments the original greedy approach with an additional exploration step guided by model kinship (highlighted in blue in Algorithm 1).
+This approach aims to merge the best-performing model with the model that has the most distinct kinship, in order to discover potentially better solutions.
+In Figure 5 (b), models generated by our strategy are marked in purple, while the best-performing models are marked in yellow.
+5.2 Results and Discussion
+Figure 6: Weight Change between two Evolution Paths.
+Figure 5 (a) illustrates the improvements in top average task performance across merging generations.
+Table 3 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
+While both strategies achieve the multitask objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72.
+In contrast, strategy utilizing model kinship escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
 Merging Models with Low Kinship can Boost Exploration.
-Figure 6(b) highlights the key branch of the model family tree.
-To investigate how merging models with low kinship helps escape local optima, we focus on the bifurcation point and analyze the weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3) in two separate experiments.
-The previous weight change, $v_{pre}$ (from Model-1-3 to Model-2-1), serves as a baseline.
-Figure 7 reveals that merging with the exploration model resulted in significant weight changes in a distinct direction, introducing novel variations into the weight space.
-In contrast, $v_{1}$ shows minimal weight change, as the merging effect is reduced due to the high similarity between the weights of openchat-3.5 and Model-2-1.
+Figure 5 (b) shows the main branch of the model family tree.
+To explore how low-kinship merges help escape local optima during saturation, we analyze weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3), with $v_{pre}$ (from Model-1-3 to Model-2-1) as baseline.
+Figure 6 shows that merging with the exploration model ($v_{2}$) induces significant, novel weight changes, while $v_{1}$ shows minimal change due to high similarity between openchat-3.5 and Model-2-1.
 Early Stopping at High Kinship can Improve Efficiency.
-We observe that the saturation stage of model evolution is particularly resource-intensive.
-In community experiments, 5 out of 14 merges in evolution path 1 resulted in an average improvement of just 0.57, while 3 out of 12 merges in evolution Path 2 yields an average improvement of 0.36.
-In our own experiments, applying a greedy strategy to a simple task lead to saturation after 2 out of 4 merges, with no further gains.
-These results indicate that human judgment and conventional stopping conditions cannot effectively halt the merging process at the optimal time.
-Therefore, we propose that model kinship can be used as an effective early stopping signal.
-When merging converges, the model kinship between top-performing models often exceeds 0.9.
-By halting the merging process at this point, time efficiency improves by approximately 30%, with minimal or no reduction in performance.
+Iterative merging is resource-intensive.
+In our main experiments, a greedy strategy saturated after 2/4 merges with no further gains.
+Looking back at community experiments, 5/14 merges in path 1 averaged only 0.57 improvement, and 3/12 merges in path 2 averaged 0.36.
+A high kinship score (PCC > 0.9) among top models may indicate convergence.
+Stopping merges early at high kinship generation saves  30% time with negligible performance loss.
 6 Related work
-Weight averaging is one of the most widely used techniques in model merging, with its origins traced back to Utans (1996), who first applied it in neural networks to achieve performance comparable to ensemble methods.
-Since the 2010s, weight averaging has found numerous applications in deep neural networks, including combining checkpoints to enhance the training process (Nagarajan & Kolter, 2019; Tarvainen & Valpola, 2017; Izmailov et al., 2018; Li et al., 2023b; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023a; Smith & Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022).
+Weight averaging, a widely used technique in model merging, traces its origins to Utans (1996).
+Since the 2010s, weight averaging has been widely applied in deep neural networks, notably for combining checkpoints to improve training stability and performance.
+(Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Smith and Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022).
 Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022).
-Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena & Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024).
+Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024).
 Ilharco et al.
 (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base.
-They demonstrate that arithmetic operations on these vectors enable model editing, such as achieving multitask learning.
 Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights.
-The Model Breadcrumbs (Davari & Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging.
+The Model Breadcrumbs (Davari and Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging.
 Merging models with different initializations requires additional considerations.
-Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh & Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024).
-While weight averaging cannot be directly applied to models with different architectures, it can still be used to enhance feasible fusion methods.
-Recent work, such as FuseChat (Wan et al., 2024b), combines weight averaging with Knowledge Fusion (Wan et al., 2024a) to develop innovative fusion techniques.
+Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh and Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024).
+While weight averaging cannot be applied to models with different architectures, it can still be used to enhance feasible fusion methods.
+Recent work like FuseChat (Wan et al., 2024b) integrates it with Knowledge Fusion (Wan et al., 2024a) to enable novel fusion approaches.
 Recently, there have been some works exploring “model evolution”.
@@
 (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models.
-Labonne (2024) develop a tool to automatically merge models on Hugging Face, using an “Automerge” experiment to explore metrics in the merging process.
+Labonne (2024) develop a tool to automatically merge models on HuggingFace.
 Akiba et al.
-(2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations, arguing that human intuition alone cannot uncover hidden patterns in merging.
-6 Conclusion and Limitations
-In this paper, we introduce model kinship, the degree of similarity or relatedness between LLMs, for merging LLMs, which can help guide our selection of candidate models.
-We conduct comprehensive experiments to demonstrate its effectiveness in understanding the model evolution process.
-We further propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship.
-We show that model kinship plays a crucial role in model evolution by guiding the process to escape local optima traps (in saturation stage), enabling further improvements.
-Additionally, we demonstrate that model kinship can detect the onset of convergence, allowing for early stopping and reducing the waste of computational resources in the merging process.
+(2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations.
+7 Conclusion
+We propose continual model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance.
+To support this framework, we introduce model kinship, a metric that guides merge candidate selection and explains both performance gains and stagnation during merging.
+Building on this, we propose Top-$k$ Greedy Merging with Model Kinship, a strategy that uses kinship to escape local optima and achieve further improvements.
+Kinship also serves as an early stopping signal by detecting convergence and reducing redundant computation.
 In a broad sense, our work explores how models can achieve autonomous evolution through model merging.
@@
 Limitations
-However, there are several limitations to consider: a) The experiments in this study are conducted exclusively on models with the Mistral architecture, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu & Dao, 2023).
-b) The analysis relies on open-source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias.
-c) Correlation metrics for model kinship have not been fully explored.
+However, there are several limitations to consider: a) The experiments in this study are conducted on models with two architectures, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu and Dao, 2023).
+Furthermore, scaling of tasks and candidate models requires further experimentation to understand the computational cost across various scenarios.
+b) The analysis is based on open source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias.
+c) Possible correlation metrics for model kinship have not been fully explored.
 Other metrics may perform better than those discussed in this paper.
 d) The effectiveness of model kinship is demonstrated through empirical evidence.
-However, a theoretical framework (such as the assumptions in Appendix C) is needed to explain model evolution and model kinship more rigorously.
-e) Model kinship currently guides merging and improves performance limits but does not support sustained evolution.
+However, a theoretical framework (such as the assumptions in Appendix 15) is needed to explain model evolution and model kinship more rigorously.
+e) Model kinship currently guides merging and enhances performance, but does not support sustained evolution.
 Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
-Reproducibility Statement
-The experimental setup can be found in Section 4.1. All model checkpoints are available on HuggingFace, with detailed information provided in Appendices A and B.
-Appendix A Details of Experiments
-All merged models from these experiments are accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets. The following tables cover two primary aspects:
+Appendix A Details of Experiments in Main Sections
+This section provides comprehensive details on the models used in the analysis of community experiments.
+The open merged models from these experiments are accessible through the Hugging Face Hub [^3] https://huggingface.co/datasets.
+The evaluation results can be accessed in the Openleaderboard [^4] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboards The following tables cover two primary aspects:
   • (1) Information on the selected model family trees for two distinct evolution paths, along with detailed analysis results for each merge.
   • (2) A summary of the merge experiments conducted for distribution analysis.
+A.1 Selecting the Evolution Path
 The evolution paths are selected using a structured process, focusing on identifying key sequences within the model family trees. The steps were as follows:
@@
   • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multitask capabilities and model kinship metrics.
-Table 3 and 4 present detailed information on the sequential merging process.
+Table 4 and 5 present detailed information on the sequential merging process.
 The second and third columns record the foundational models involved in each merge, while the final column lists the resulting merged models.
-Table 3: Model Family tree of evolution Path 1.
+A.2 Additional Results in Analysis
+Table 6 and Table 7 present detailed analysis results that are not reported in the main paper.
+These include Average Task Performance (ATP), merge gains, and model kinship values, which are computed using Pearson Correlation coefficient, Cosine Similarity, and Euclidean Distance for each merge.
+Table 8 presents all merge experiments contributing to the distribution analysis.
+The selection of sample experiments adheres to two rules: (1) Samples are evenly chosen across average task performance values ranging from 0.7 to 0.7686 (the average task performance of the best 7B merged model) to accurately reflect the full scope of model evolution.
+(2) The experiments involve merges of two foundation models, as including multiple models introduces excessive noise.
+A.3 Details of Datasets Selection
+In the main experiments, we utilize three task-specific benchmark datasets—Winogrande, GSM8k, and TruthfulQA—to evaluate the distinct strengths of the models.
+These datasets assess the following capabilities:
+  • Winogrande: Evaluates the model’s commonsense reasoning.
+  • GSM8k: Measures the model’s mathematical reasoning.
+  • TruthfulQA: Assesses the model’s ability to identify and address human falsehoods.
 Gen Model-1 Model-2 Model-Merged
@@
 14 YamShadow-7B Experiment28 YamshadowExperiment28-7B
-Table 4: Model Family tree of evolution Path 2.
+Table 4: Model Family tree of evolution Path 1.
 Gen Model-1 Model-2 Model-Merged
@@
 12 Experiment28-7B YamShadow-7B YamshadowExperiment28-7B
-Table 5 and Table 6 present detailed analysis results that are not reported in the main paper.
-These include Average Task Performance (ATP), merge gains, and model kinship values, which are computed using Pearson Correlation coefficient, Cosine Similarity, and Euclidean Distance for each merge.
-Table 5: Summary of Path 1 Results.
+Table 5: Model Family tree of evolution Path 2.
 Gen Model-Merged ATP Gain PCC CS ED
@@
 14 YamshadowExperiment28-7B 76.86 0.25 0.98 0.98 1.61
-Table 6: Summary of Path 2 Results.
+Table 6: Summary of Path 1 Results.
 Gen Model-Merged ATP Gain PCC CS ED
@@
 12 YamshadowExperiment28-7B 76.86 0.25 0.98 0.98 1.61
-Table 7 presents all merge experiments contributing to the distribution analysis.
-The selection of sample experiments adheres to two rules: (1) Samples are evenly chosen across average task performance values ranging from 0.7 to 0.7686 (the average task performance of the best 7B merged model) to accurately reflect the full scope of model evolution.
-(2) The experiments involve merges of two foundation models, as including multiple models introduces excessive noise.
-Table 7: All Sample Experiments used in distribution analysis.
+Table 7: Summary of Path 2 Results.
 Model 1 Model 2 Merge Gain
@@
 YugoGPT AlphaMonarch-7B -5.96
-Appendix B Details of Model Group Selection
-This appendix presents the exact models included in each model group, as shown in Table 8.
+Table 8: All Sample Experiments used in distribution analysis.
+Appendix B Full Evaluation Results of Main Experiments and Additional Experiments
+B.1 Main Mistral-7B Experiments
+Table 9 provides a detailed evaluation of the main experiments, including the results for the exploration models and their performance on specific tasks.
+The model kinship experiment is terminated early at generation 5, as a more promising evolution path is subsequently identified.
+B.2 Additional Experiments
+To assess the generalization of our strategy across different model architectures and task sets, we conduct two additional experiments.
+B.2.1 Mistral-7B Experiments with a Different Task Set
+We perform further evaluations using Mistral-7B, based on three distinct foundation models: MistralHermes-CodePro-7B-v1, metamath-mistral-7b, and open-chat-3.5-1210.
+These models are assessed on the HumanEval, GSM8k, and TruthfulQA benchmarks.
+The model kinship-based merging process is terminated early at generation 3, as a more promising evolution trajectory is identified.
+In this task setting, model kinship-guided exploration successfully discovers models (e.g., Child3-3) that significantly outperform their respective initial baselines.
+B.2.2 LLaMA-2-8B Experiments
+We further evaluate our strategy on LLaMA-2-8B using three task-specific fine-tuned models.
+Table 11 summarizes the results of these additional experiments.
+The model kinship-based merging process is terminated early at generation 6 upon the discovery of a more favorable evolutionary path.
+Consistent with the results from Mistral-7B, model evolution guided by model kinship continues to facilitate performance improvements beyond the capabilities of the original models.
+Model TruthfulQA Winogrande GSM8K Avg. Model Kinship
+MetaMath 44.89 75.77 70.51 63.72 /
+Instruct 68.26 77.19 40.03 61.82 /
+Open-chat 52.15 80.74 65.96 66.28 /
+model-1-1-greedy 52.51 76.16 57.85 62.17 0.01
+model-1-2-greedy 58.04 76.32 57.72 64.02 -0.02
+model-1-3-greedy 48.96 78.69 72.86 66.84 0.05
+model-2-1-greedy 50.94 80.11 75.13 68.72 0.93
+model-2-2-greedy 49.78 78.93 55.72 61.47 0.57
+model-2-3-greedy 52.36 78.61 52.99 61.32 0.58
+model-2-exp 61.01 79.56 63.76 68.11 -0.02
+model-3-1-greedy 51.95 80.51 73.31 68.59 0.95
+model-3-2-greedy 49.96 79.72 73.54 67.74 0.93
+model-3-3 56.95 80.25 70.00 69.06 0.24
+model-3-4 54.38 78.45 73.01 68.61 0.32
+model-3-exp 54.13 78.69 71.65 68.15 0.03
+model-4-1-greedy 50.82 80.11 74.60 68.51 0.98
+model-4-2-greedy 50.36 79.47 74.31 68.04 0.98
+model-4-3-greedy 51.04 79.72 74.83 68.53 0.94
+model-4-4 53.31 79.40 73.54 68.75 0.54
+model-4-5 52.48 79.01 73.68 68.39 0.66
+model-4-6 53.69 79.72 73.69 69.03 0.52
+model-4-exp 55.16 78.53 71.80 68.49 0.48
+model-5-1 54.85 79.37 73.31 69.13 0.65
+model-5-2 54.78 79.40 72.86 68.98 0.65
+model-5-3 53.49 79.24 73.16 68.63 0.98
+model-5-exp 52.98 79.32 72.78 68.36 0.59
+Table 9: Evaluation Results of Main Experiments of Mistral-7B.
+Model TruthfulQA GSM8K HumanEval Avg. Model Kinship
+MetaMath 44.89 70.51 17.68 44.36 /
+Openchat-3.5-1210 52.15 65.96 2.44 40.18 /
+MistralHermes-CodePro-7B-v1 49.67 60.88 22.56 44.37 /
+child1-1-greedy 51.87 69.60 15.80 45.76 0.19
+child1-2-greedy 48.04 72.78 9.15 43.32 0.08
+child1-3-greedy 48.96 72.86 18.29 46.70 0.05
+child2-1-greedy 50.24 71.72 12.20 44.72 0.15
+child2-2-greedy 50.88 73.24 7.32 43.81 0.92
+child2-3-greedy 51.15 67.32 19.51 45.99 0.34
+child2-4-greedy 50.05 72.33 4.88 42.42 0.21
+child2-exp 50.33 71.11 18.90 46.78 0.21
+child3-1-greedy 51.47 69.22 21.34 47.34 0.73
+child3-2-greedy 50.71 72.40 9.15 44.09 0.82
+child3-3 49.69 74.37 21.34 48.47 0.82
+child3-4 50.57 69.75 17.68 46.00 0.91
+child4-1-greedy 50.56 68.46 12.20 43.74 0.79
+child4-2-greedy 51.28 68.46 19.51 46.42 0.95
+child5-1-greedy 51.36 68.69 20.73 46.93 0.99
+child5-2-greedy 50.49 73.24 9.76 44.50 0.78
+child6-1-greedy 50.50 73.24 9.15 44.30 0.78
+child6-2-greedy 51.42 69.14 20.12 46.89 0.99
+child7-1-greedy 51.36 68.82 20.34 46.84 0.99
+child7-2-greedy 51.42 68.74 20.81 46.99 0.99
+child7-3-greedy 51.44 69.15 20.44 47.01 0.99
+Table 10: Evaluation Results of Additional Experiments of Mistral-7B.
+Model TruthfulQA Winogrande GSM8K Avg. Model Kinship
+winogrande 42.0 77.9 6.4 42.1 /
+GSM8K 39.0 73.4 38.0 50.1 /
+TruthfulQA 56.7 68.9 9.5 45.0 /
+child1-1-greedy 40.2 79.3 34.2 51.2 0.03
+child1-2-greedy 46.7 74.4 34.2 51.7 0.01
+child1-3-greedy 46.1 77.1 1.9 41.7 0.01
+child-2-1-greedy 44.5 78.5 36.8 53.2 0.19
+child-2-2-greedy 43.7 74.0 40.4 52.7 0.45
+child-2-3-greedy 38.9 77.5 37.1 51.1 0.39
+child-2-exp 43.3 81.2 28.5 51.0 0.01
+child-3-1-greedy 44.2 77.1 37.3 52.8 0.88
+child-3-2-greedy 45.4 77.5 34.5 52.4 0.79
+child-3-3-greedy 45.0 73.8 36.6 51.8 0.89
+child-3-exp 45.1 78.6 30.3 51.3 0.58
+child-4-1-greedy 44.4 78.5 36.8 53.2 0.95
+child-4-2-greedy 44.1 75.5 40.0 53.1 0.97
+child-4-exp 43.3 80.9 32.6 52.2 0.81
+child-5-1-greedy 44.2 77.1 37.2 52.8 0.97
+child-5-2-greedy 44.3 77.4 36.7 52.8 0.91
+child-5-3-greedy 44.3 78.3 36.8 53.1 0.98
+child-5-exp 44.5 78.1 32.0 51.5 0.64
+child-6-1-greedy 44.5 78.5 36.8 53.2 0.99
+child-6-2-greedy 44.4 78.3 36.8 53.2 0.99
+child-6-3-greedy 44.3 78.3 36.8 53.1 0.99
+child-6-exp 44.3 80.4 35.3 53.4 0.80
+Table 11: Evaluation Results of addtional experiments of Llama-2.
+Appendix C Algorithm Details for the Main Experiment
+In this section, we present the detailed algorithms employed in our main experiment, along with an ablation study to validate our baseline method, Top $k$ Greedy Merging.
+C.1 Algorithms
+The Top-$k$ Greedy Merging algorithm aims to iteratively construct improved models through pairwise merging, guided by performance evaluation and, in the extended version, model kinship.
+It begins with a set of $n$ foundation models $M=\{m_{1},m_{2},\dots,m_{n}\}$.
+In the first step, all possible pairs of models are merged to form the first generation $G_{1}$ of merged models.
+These new models are added back into the candidate set $M$.
+The algorithm then evaluates all models in $M$ using a task-specific evaluation function $f$ and selects the top $k$ performing models to form the working set $S$.
+It maintains an iterative loop that continues as long as the top-$k$ set $S$ changes between iterations, ensuring exploration continues only while performance improves.
+Within each iteration, all model pairs from $S$ are merged to produce the next generation of models $G_{i}$.
+These new models are added into $M$, and their performance is evaluated using $f$ to update $S$.
+In the variant with model kinship, additional steps introduce a model exploration mechanism.
+This kinship-guided exploration step is designed to escape local optima by encouraging diversity in the merging path, potentially leading to models with better generalization or complementary capabilities.
+The algorithm terminates when the top-$k$ set $S$ stabilizes, indicating no further performance gains are observed.
+Each model is named according to its generation and creation order to track its origin during analysis.
+C.2 Ablation Study of Greedy Strategy
+The ablation study on the Greedy Strategy is conducted using the Mistral-7B architecture, following the same experimental settings outlined in the main experiments.
+For comparison, we employ the random-merge strategy, where models in each generation are merged with randomly selected models (excluding themselves) from the repository, as illustrated in Algorithm 2.
+The following table presents the evaluation results. Each column represents:
+  • Model: The name of each model. Note that the first three entries are fine-tuned foundation models used in our experiments.
+  • TruthfulQA_mc2, Winogrande, GSM8K: The benchmark results for each dataset, indicating the model’s task-specific capabilities.
+  • Average: The average score across all benchmarks, reflecting the model’s overall generalization performance.
+  • Model Kinship: The kinship score (Here, we use cosine similarity to measure model kinship) of the parent models involved in the merge, indicating their relatedness.
+  • Parent-1 and Parent-2: The names of the parent models used in the merging process.
+In the random-merge strategy, the average performance in each generation fluctuates.
+The highest average performance achieved is 68.55, slightly lower than the 68.72 observed in the greedy experiment.
+While the random-merge strategy avoids convergence to local optima, it demonstrates an unstable improvement process, which can lead to unpredictable results.
+Model TruthfulQA_mc2 Winogrande GSM8K Average Model Kinship
+MetaMath-mistral-7B 44.89 75.77 70.51 63.72 /
+Mistral-7B-Instruct-v0.2 68.26 77.19 40.03 61.82 /
+Open-chat-3.5-1210 52.15 80.74 65.96 66.28 /
+child1-1 52.51 76.16 57.85 62.17 0.01
+child1-2 58.04 76.32 57.72 64.02 0.01
+child1-3 48.96 78.69 72.86 66.84 0.03
+child2-1 44.68 74.00 50.80 56.40 0.29
+child2-2 49.78 78.93 55.72 61.47 0.41
+child2-3 61.01 79.56 63.76 68.11 0.01
+child3-1 51.52 78.23 56.71 62.15 0.84
+child3-2 43.52 75.22 47.43 55.39 0.59
+child3-3 54.32 78.53 72.81 68.55 0.28
+child4-1 55.32 78.41 56.23 63.32 0.54
+child4-2 50.53 78.42 57.65 62.20 0.86
+child4-3 53.45 79.31 72.65 68.47 0.67
+Table 12: Evaluation results using the random-merge strategy.
+Algorithm 1 Top $k$ Greedy Merging with Model Kinship.
+0:  A set $M$ of $n$ foundation models $\{m_{1},m_{2},\dots,m_{n}\}$, Evaluation function $f$, Similarity metric function $sim(\cdot,\cdot)$ for model kinship.
+1:  Generate the first generation of merged models $G_{1}$ by merging each pair in set $M$, and set generation $i$ = 1.
+2:  Combine the set $G_{1}$ into set $M$.
+3:  Evaluate each model $m$ in set $M$, and select the top $k$ models. Denote this set as $S=\{m_{1},m_{2},\dots,m_{k}\}$.
+4:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $k$ models from the previous iteration.
+5:  while $S\neq S_{\text{prev}}$ do
+6:     i++
+7:     Set $S_{\text{prev}}=S$.
+8:     Select each model pair from $S$. Denote this set as $P=\{p_{1},p_{2},\dots,p_{j}\}$.
+9:     Merge every selected pair in set $P$ as merged model set $G_{i}=\{m_{1},m_{2},\dots,m_{j}\}$ for generation $i$, and add each merged model into set $M$.
+10:     Identify the current best model $m_{best}\in S$.
+11:     Identify the model $m_{f}\in S$ with the lowest model kinship to $m_{best}$ from the $G_{i-1}$ according to the similarity metric $sim(\cdot,\cdot)$.
+12:     Merge $m_{f}$ with $m_{best}$ to generate a new model $m_{\text{exp}}$, and add $m_{\text{exp}}$ into set $G_{i}$ and set $M$.
+13:     Evaluate each new model $m\in G_{i}$ using $f$ and update $S$.
+14:     Evaluate $m_{\text{exp}}$ using $f$ and update $S$.
+15:  end while
+Note: The blue-highlighted steps are only executed in modified experiments incorporating model kinship-based exploration.
+To distinguish between different models in the subsequent experiments, each model generated in a given generation is named as model-generation-id.
+Algorithm 2 Random Merge Algorithm.
+0:  A set $M$ of $n$ foundation models $\{m_{1},m_{2},\dots,m_{n}\}$, Evaluation function $f$.
+1:  Generate the first generation of merged models $G_{1}$ by randomly merging pairs in set $M$, and set generation $i$ = 1.
+2:  Combine the set $G_{1}$ into set $M$.
+3:  Evaluate each model $m$ in set $M$.
+4:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $k$ models from the previous iteration.
+5:  while $S\neq S_{\text{prev}}$ do
+6:     i++
+7:     Set $S_{\text{prev}}=S$.
+8:     Randomly select pairs of models from $M$. Denote this set as $P=\{p_{1},p_{2},\dots,p_{j}\}$.
+9:     Merge each selected pair in set $P$ to form the merged model set $G_{i}=\{m_{1},m_{2},\dots,m_{j}\}$ for generation $i$, and add each merged model into set $M$.
+10:     Evaluate each new model $m\in G_{i}$ using $f$ and update $S$.
+11:  end while
+Appendix D Additional Analysis for Community Model Evolution
+D.1 Analysis of Model Kinship Change across Merging Stages
+To determine whether the discovery of increasing model kinship in model evolution paths can be generalized to the entire model evolution process, we perform an analysis of the merging stages.
+Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of models at each stage.
+For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
+D.2 Details of Model Group Selection
+This section presents the exact models included in each model group, as shown in Table 13.
 The selection process is conducted across three distinct groups: (1) the top 5 models on the leaderboard, with a performance difference of 0.2, (2) 5 models with performance scores around 73 and a performance difference of 0.2, and (3) 5 fine-tuned models.
-It is important to note that the fine-tuned models were not selected based on performance, and may exhibit significant differences in results.
-Table 8: Model Group in Kinship Matrix Analysis.
+It is important to note that the fine-tuned models are not selected based on performance, and may exhibit significant differences in results.
 Group Models
@@
 WizardLM-2
-Appendix C Assumption of Continual Model Merging
-Figure 8: An intuitive illustration of the optimization process assumption in model evolution, where models progressively converge towards the optimal model.
-Our findings in the main paper offer a new perspective on model evolution through multiple merging.
+Table 13: Model Group in Kinship Matrix Analysis.
+Figure 7 illustrates the model kinship between models within each group.
+We observe that model kinship increases with the average task performance across models that follow different evolution paths.
+Additionally, during the saturation stage, all potential merges display a strong affinity, with model kinship values nearing 1.
+Figure 7: The Model Kinship Matrices for the three model groups.
+Each element represents the model kinship value between the corresponding models.
+In Group B and C, the merged models are arranged by average task performance, ordered from high to low (left to right).
+Appendix E Analysis between Task Relatedness and Model Kinship
+In the formulation of model kinship, we use the placeholder (,$\sim$($\cdot$,$\cdot$)∼ ( ⋅ , ⋅ ) as a similarity metric function to explore options that can effectively capture task-related differences.
+One such metric is cosine similarity, derived from the analysis in the task vector, which has been validated as effective for representing differences in single-task models through the cosine similarity of delta parameters (task vectors).
+In addition to cosine similarity, we also investigate the Pearson correlation coefficient and Euclidean distance.
+However, we have not thoroughly evaluated the applicability of these metrics in the context of model evolution, particularly for merged models with multitask capabilities.
+To address this, we examine the relationship between the similarity metrics and task information in subsequent sections.
+Our analysis focuses on the LLaMA-2 architecture, as we can find the necessary open-source fine-tuned checkpoints on various datasets.
+To measure differences between models, we currently use a preliminary evaluation method: the Average Task Performance Difference (ATPD), which aims to represent task capability differences based on evaluation performance.
+The Average Task Performance Difference (ATPD) between two models, $M_{1}$ and $M_{2}$, is calculated by averaging the absolute differences in performance across all tasks.
+Let $T$ denote the set of tasks, and $P_{i}^{(j)}$ represent the performance of model $M_{j}$ on task $i$.
+Then, the ATPD is defined as:
+$\text{ATPD}(M_{1},M_{2})=\frac{1}{|T|}\sum_{i\in T}\left|P_{i}^{(1)}-P_{i}^{(2)}\right|$
+  • $|T|$: the total number of tasks.
+  • $P_{i}^{(1)}$ and $P_{i}^{(2)}$: performances of models $M_{1}$ and $M_{2}$ on task $i$.
+  • $\left|P_{i}^{(1)}-P_{i}^{(2)}\right|$: absolute difference in performance for task $i$.
+Method Corr(cs) Corr(pcc) Corr(ed)
+Value -0.77 -0.74 0.80
+Table 14: Correlation values between ATPD and model kinship.
+For this study, we utilize models from additional LLaMA-2 experiments (Appendix.B).
+These models are merged from three fine-tuned models, allowing us to control the generated models to focus solely on the corresponding task capabilities.
+The following table presents the results, with Winogrande, TruthfulQA, and GSM8K representing the performance differences across each task.
+The results in Table.14 demonstrate strong correlations: Cosine Similarity (-0.77) and Pearson Correlation Coefficient (-0.74) exhibit negative correlations, while Euclidean Distance (0.80) shows a positive correlation.
+This supports that model kinship is related to task differences.
+As mentioned in the limitations, the current metrics are viable but not optimal.
+Combining them with task information studies could further enhance the value of our work.
+E.1 Additional Results: Analysis of Model Kinship and Average Task Performance
+This section provides supplementary analysis on the relationship between model kinship and average task performance.
+Figure 8 illustrates a comparison between average task performance and model kinship using two additional metrics not included in the main paper.
+From an intuitive observation, model kinship based on the three metrics exhibits a similar correlation with average task performance.
+Model 1 Model 2 Winogrande TruthfulQA GSM8K ATPD Kinship(cs) Kinship(pcc) Kinship(ed)
+child-4-1-greedy child-5-3-greedy 0.10 0.00 0.20 0.10 0.99 0.99 2.17
+child-2-1-greedy child-4-1-greedy 0.20 0.10 0.00 0.10 0.98 0.99 4.22
+child-2-1-greedy child-5-3-greedy 0.10 0.10 0.20 0.13 0.99 0.99 2.19
+child-4-exp child-2-1-greedy 1.10 0.90 0.10 0.70 0.80 0.75 25.53
+child-2-1-greedy child-3-1-greedy 0.20 1.30 0.70 0.73 0.95 0.98 6.74
+child-4-1-greedy child-6-exp 0.10 1.90 1.40 1.13 0.74 0.71 25.54
+child-4-1-greedy child-4-2-greedy 0.30 3.00 3.20 2.17 0.97 0.98 6.57
+child-2-2-greedy child-3-1-greedy 0.50 3.10 3.10 2.23 0.97 0.98 6.57
+child-2-1-greedy child-4-2-greedy 0.50 3.10 3.20 2.27 0.91 0.96 9.29
+child-3-exp child-2-1-greedy 0.70 0.20 6.30 2.40 0.64 0.52 35.52
+child-4-exp child-2-1-greedy 1.10 2.50 4.00 2.53 0.78 0.75 25.53
+child-2-1-greedy child1-2-greedy 2.30 4.00 2.40 2.90 0.79 0.89 15.75
+child-2-1-greedy child-2-2-greedy 0.70 4.40 3.80 2.97 0.88 0.95 12.43
+child-2-2-greedy child1-2-greedy 3.00 0.40 6.20 3.20 0.89 0.92 11.68
+child1-1-greedy GSM8K 1.20 5.90 3.80 3.63 0.39 0.46 36.39
+child1-1-greedy child1-2-greedy 6.50 4.90 0.00 3.80 0.19 0.16 38.07
+child-2-exp child-2-1-greedy 1.10 2.80 8.10 4.00 0.58 0.77 28.33
+child1-2-greedy GSM8K 7.70 1.00 3.80 4.17 0.45 0.38 26.32
+child-2-1-greedy child1-3-greedy 7.80 3.10 2.90 4.60 0.58 0.51 45.24
+child-3-1-greedy child-2-exp 0.90 4.10 8.80 4.60 0.58 0.63 32.45
+winogrande TruthfulQA 14.70 9.00 3.10 8.93 0.01 0.01 74.49
+child1-2-greedy child1-3-greedy 0.60 2.70 32.30 11.87 0.64 0.52 46.06
+child1-2-greedy winogrande 4.70 3.50 27.80 12.00 0.01 0.02 55.89
+winogrande GSM8K 3.00 4.50 31.60 13.03 0.03 0.11 54.01
+child1-1-greedy child1-3-greedy 5.90 2.20 32.30 13.47 0.52 0.64 44.16
+GSM8K TruthfulQA 17.70 4.50 28.50 16.90 0.01 0.01 61.56
+Table 15: Summary of Model Merging Results.
+Figure 8: Illustration of comparison between the correlation of Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED) with average task performance (Normalized to the same value scale).
+Appendix F Optimization Assumption of Model Evolution
+Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy.
+It shows that Merged Model 2 may be overlooked, despite its potential for better multitask performance.
+Our findings using new strategy offer a new perspective on model evolution through multiple merging.
 If the merging process can be improved using a common optimization strategy, it raises the question of whether the underlying mechanism mirrors this optimization problem.
@@
 Hypothesis: The evolution process may be simplified to a binary search process for most weight-averaging-based model merging methods.
-Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy.
-It shows that Merged Model 2 may be overlooked, despite its potential for better multitask performance.
 Figure 10 illustrates the ideal scenario in our assumption where multiple merges produce a highly generalized model.
@@
 However, we currently lack sufficient evidence to validate this hypothesis. Future work is needed to explore this further.
-Appendix D Other Information and Results
-This section presents additional results that are not reported in the main paper but may be useful for those interested.
-Figure 10 compares average task performance and model kinship using alternative metrics.
-Table 9 provides information on the exploration models used in our experiments.
-Figure 11 depicts a more “realistic” representation of combining two features/capabilities in evolution, highlighting the challenges in the selection process.
-Figure 10: Illustration of comparison between the correlation of Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED) with average task performance (Normalized to the same value scale).
-Table 9: Results of Exploration Models in our modified strategy.
-Model Avg. Gain Kinship
-model-2-exp 68.11 +3.78 -0.02
-model-3-exp 68.15 +2.70 0.03
-model-4-exp 68.49 +3.42 0.48
-model-5-exp 68.36 -0.09 0.53
+Figure 10: An intuitive illustration of the optimization process assumption in model evolution, where models progressively converge towards the optimal model.
 Figure 11: An intuitive comparison between selective breeding and continual model merging.
@@
 The right process shows developing a model with capabilities of coding and math through model evolution.
+Appendix G Referenced Concepts in Evolutionary biology
+In this section, we detail the conceptual parallels between biological processes and model merging, highlighting our motivation for employing model kinship.
+G.1 Iterative Merging vs. Artificial Selection
+We draw inspiration for model evolution from biological evolution, specifically focusing on the correlation between biological evolution through artificial selection and model evolution via model merging.
+Artificial selection involves retaining desirable traits by manually selecting breeding pairs in each generation, typically those exhibiting the most significant features.
+Similarly, model evolution, as explored in this paper through Iterative Model Merging, adopts a comparable approach: users preserve desired task capabilities by strategically selecting merging pairs.
+Through iterative merging, they can develop a model that is proficient in all tasks in a given task set.
+To illustrate this comparison more effectively, Figure 11 shows an example of combining two features/task capabilities in evolution.
+G.2 Inbreeding Depression vs. Saturation Stage
+As highlighted in the main paper, one of our key findings is that the late stage of model evolution often enters a saturation stage, during which models exhibit minimal differences from one another.
+This phenomenon parallels "inbreeding depression" in artificial selection, where breeding closely related individuals reduces genetic diversity and fitness.
+Although genetic inheritance and model weights operate differently, merging closely related models leads to new models with minimal variation, thereby reducing the effectiveness of merging, particularly in weight averaging.
+To address this issue, we propose quantifying the differences between models, a concept we term model kinship, to guide the merging process and mitigate the challenges associated with the saturation stage in model evolution.
+Appendix H References to Open Models
+See Table LABEL:tab:ref.
+Table 16: Model and Hugging Face Reference Links
+Model Name Hugging Face Reference
+Multi_verse_model-7B https://huggingface.co/MTSAIR/multi_verse_model
+Experiment26-7B https://huggingface.co/yam-peleg/Experiment26-7B
+M7-7b https://huggingface.co/liminerity/M7-7b
+StrangeMerges_32-7B-slerp https://huggingface.co/Gille/StrangeMerges_32-7B-slerp
+Ognoexperiment27 https://huggingface.co/automerger/OgnoExperiment27-7B
+YamShadow-7B https://huggingface.co/automerger/YamShadow-7B
+Experiment28 https://huggingface.co/yam-peleg/Experiment28-7B
+shadow-clown-7B-slerp https://huggingface.co/CorticalStack/shadow-clown-7B-slerp
+yam-jom-7B https://huggingface.co/mayacinka/yam-jom-7B
+StrangeMerges_21-7B-slerp https://huggingface.co/Gille/StrangeMerges_21-7B-slerp
+StrangeMerges_31-7B-slerp https://huggingface.co/Gille/StrangeMerges_31-7B-slerp
+NeuralBeagle14-7B https://huggingface.co/mlabonne/NeuralBeagle14-7B
+Turdus https://huggingface.co/udkai/Turdus
+DareBeagle-7B https://huggingface.co/shadowml/DareBeagle-7B
+TurdusBeagle-7B https://huggingface.co/leveldevai/TurdusBeagle-7B
+FernandoGPT-v1 https://huggingface.co/samir-fama/FernandoGPT-v1
+StrangeMerges_10-7B-slerp https://huggingface.co/Gille/StrangeMerges_10-7B-slerp
+TurdusDareBeagle-7B https://huggingface.co/leveldevai/TurdusDareBeagle-7B
+MarcMistral-7B https://huggingface.co/flemmingmiguel/MarcMistral-7B
+StrangeMerges_20-7B-slerp https://huggingface.co/Gille/StrangeMerges_20-7B-slerp
+NeuTrixOmniBe-7B-model-remix https://huggingface.co/Kukedlc/NeuTrixOmniBe-7B-model-remix
+StrangeMerges_11-7B-slerp https://huggingface.co/Gille/StrangeMerges_11-7B-slerp
+MBX-7B-v3 https://huggingface.co/flemmingmiguel/MBX-7B-v3
+Marcoroni-7B-v3 https://huggingface.co/AIDC-ai-business/Marcoroni-7B-v3
+Mistral-7B-Merge-14-v0.1 https://huggingface.co/EmbeddedLLM/Mistral-7B-Merge-14-v0.1
+distilabeled-Marcoro14-7B-slerp https://huggingface.co/argilla/distilabeled-Marcoro14-7B-slerp
+UNA-TheBeagle-7b-v1 https://huggingface.co/fblgit/UNA-TheBeagle-7b-v1
+CatPPT-base https://huggingface.co/rishiraj/CatPPT-base
+CatMacaroni-Slerp https://huggingface.co/cookinai/CatMacaroni-Slerp
+LeoScorpius-7B https://huggingface.co/viethq188/LeoScorpius-7B
+NeuralDaredevil-7B https://huggingface.co/mlabonne/NeuralDaredevil-7B
+StrangeMerges_9-7B-dare_ties https://huggingface.co/Gille/StrangeMerges_9-7B-dare_ties
+mistral-ft-optimized-1218 https://huggingface.co/OpenPipe/mistral-ft-optimized-1218
+NeuralHermes-Mistral-2.5-7B https://huggingface.co/mlabonne/NeuralHermes-2.5-Mistral-7B
+neural-chat-7b-v3-2 https://huggingface.co/Intel/neural-chat-7b-v3-2
+OpenHermes-2.5-Mistral-7B https://huggingface.co/teknium/OpenHermes-2.5-Mistral-7B
+StrangeMerges_30-7B-slerp https://huggingface.co/Gille/StrangeMerges_30-7B-slerp
+Experiment24 https://huggingface.co/yam-peleg/Experiment24-7B
+neural-chat-7b-v3-3 https://huggingface.co/Intel/neural-chat-7b-v3-3
+MultiverseEx26-7B-slerp https://huggingface.co/allknowingroger/MultiverseEx26-7B-slerp
+CalmExperiment-7B-slerp https://huggingface.co/allknowingroger/CalmExperiment-7B-slerp
+CapybaraMarcoroni-7B https://huggingface.co/AtAndDev/CapybaraMarcoroni-7B
+DistilHermes-2.5-Mistral-7B https://huggingface.co/eren23/DistilHermes-2.5-Mistral-7B
+Calme-7B-Instruct-v0.9 https://huggingface.co/MaziyarPanahi/Calme-7B-Instruct-v0.9
+StrangeMerges_16-7B-slerp https://huggingface.co/Gille/StrangeMerges_16-7B-slerp
+coven_7b_128k_orpo_alpha https://huggingface.co/raidhon/coven_7b_128k_orpo_alpha
+Kunoichi-DPO-v2-7B https://huggingface.co/SanjiWatsuki/Kunoichi-DPO-v2-7B
+AlphaMonarch-7B https://huggingface.co/mlabonne/AlphaMonarch-7B
+StrangeMerges_15-7B-slerp https://huggingface.co/Gille/StrangeMerges_15-7B-slerp
+Kunoichi-7B https://huggingface.co/SanjiWatsuki/Kunoichi-7B
+Mistral-T5-7B-v1 https://huggingface.co/ignos/Mistral-T5-7B-v1
+Marcoroni-neural-chat-7B-v2 https://huggingface.co/Toten5/Marcoroni-neural-chat-7B-v2
+Marcoro14-7B-slerp https://huggingface.co/Rupesh2/Marcoro14-7B-slerp
+MarcDareBeagle-7B https://huggingface.co/leveldevai/MarcDareBeagle-7B
+MarcBeagle-7B https://huggingface.co/leveldevai/MarcBeagle-7B
+MetaMath-Mistral-7B https://huggingface.co/meta-math/MetaMath-Mistral-7B
+openchat-3.5-1210 https://huggingface.co/openchat/openchat-3.5-1210
+Tulpar-7b-v2 https://huggingface.co/HyperbeeAI/Tulpar-7b-v2
+YugoGPT https://huggingface.co/gordicaleksa/YugoGPT

### v3

@@
 Abstract
-Model merging has become one of the key technologies for enhancing the capabilities and efficiency of Large Language Models (LLMs).
-The open-source community has driven model evolution by iteratively merging existing models.
-However, a principled understanding of the expected gains and underlying factors in model merging remains lacking.
-In this work, we examine model evolution through continual merging, analogous to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs.
-With comprehensive empirical analysis, we find that there is a certain relationship between model kinship and the performance gains after model merging, which can help guide our selection of candidate models.
-Inspired by this, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can yield better performance on benchmark datasets.
-Specifically, we discover that using model kinship as a criterion can assist us in continuously performing model merging, alleviating the degradation (local optima) in model evolution, whereas model kinship can serve as a guide to escape these traps [^1] Code is available at https://github.com/zjunlp/ModelKinship..
+Model merging has emerged as a key technique for enhancing the capabilities and efficiency of Large Language Models (LLMs).
+The open-source community has driven model evolution by iteratively merging existing models, yet a principled understanding of the gains and underlying factors in model merging remains limited.
+In this work, we study model evolution through iterative merging, drawing an analogy to biological evolution, and introduce the concept of model kinship, the degree of similarity or relatedness between LLMs.
+Through comprehensive empirical analysis, we show that model kinship is closely linked to the performance improvements achieved by merging, providing a useful criterion for selecting candidate models.
+Building on this insight, we propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship, which can improve benchmark performance.
+Specifically, we discover that incorporating model kinship as a guiding criterion enables continuous merging while mitigating performance degradation caused by local optima, thereby facilitating more effective model evolution [^1] Code is available at https://github.com/zjunlp/ModelKinship..
 Exploring Model Kinship for Merging Large Language Models
-Yedi Hu, Yunzhi Yao, Shumin Deng , Huajun Chen, Ningyu Zhangthanks: Corresponding Author. Zhejiang University   National University of Singapore, NUS-NCS Joint Lab, Singapore {zhangningyu}@zju.edu.cn
+Yedi Hu, Yunzhi Yao, Ningyu Zhang, Huajun Chen, Shumin Dengthanks: Corresponding Author.
+Zhejiang University   National University of Singapore, NUS-NCS Joint Lab {231sm,zhangningyu}@zju.edu.cnshumin@nus.edu.sg
 1 Introduction
-Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, particularly demonstrating significant effectiveness in Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023).
-However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives the increasing demand for multitask learning solutions (Zhang and Yang, 2022; Lu et al., 2024; Liu et al., 2024).
+Fine-tuning pre-trained models (PTMs) for downstream tasks has become a popular practice, and has proven particularly effective for Large Language Models (LLMs) (Kolesnikov et al., 2020; Qiu et al., 2020; Askell et al., 2021; Ouyang et al., 2022; Zhao et al., 2023).
+However, deploying separate fine-tuned models for each task can be resource-intensive (Fifty et al., 2021), which drives increasing demands for multi-task learning solutions (Zhang and Yang, 2022; Yu et al., 2024a; Lu et al., 2024; Liu et al., 2024).
 Figure 1: An intuitive comparison between wheat evolution and model evolution.
-An interesting parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b).
+A parallel can be drawn between biological reproduction (Part a) and the process of model evolution (Part b).
 In biological systems, offspring inherit genetic material from both parents, forming a new genotype through the combination of parental traits.
@@
 Notably, model kinship can serve as an effective tool to guide this iterative model merging process (e.g., infer whether there may be gains after model merging.).
-Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multitask objectives by integrating multiple expert models.
-Furthermore, advancements in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) enable users with limited expertise to easily conduct merging experiments, leading to an evolution of LLMs for the community.
-To date, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023).
-Many of these models are created through a biologically inspired evolutionary process that involves iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)).
-Despite these successes, the current approach faces critical limitations.
-Progress often relies on trial and error and extensive human expertise, with little formal guidance or standardized procedure.
-To address this problem, we propose a strategic model evolution framework (Figure 1(c)) that leverages explicit strategies to guide the direction of model evolution toward improved performance.
+Recent studies suggest that model merging (Singh and Jaggi, 2020; Sung et al., 2023; Goddard et al., 2024; Matena and Raffel, 2022; Yang et al., 2024a; Jang et al., 2024) offers a viable approach for achieving multi-task objectives by integrating multiple expert models.
+Furthermore, advances in model merging toolkits (Goddard et al., 2024; Tang et al., 2024) have lowered the technical barrier, allowing users with limited expertise to easily conduct merging experiments, thus leading to an evolution of LLMs for the community.
+At present, researchers have developed various powerful LLMs using model merging techniques (Beeching et al., 2023).
+Many of these models are created through a biologically inspired evolutionary process involving iterative merging, an approach that we refer to as model evolution (Figure 1(a,b)).
+Despite these successes, current merging practice faces critical limitations.
+Progress often relies on continuous trial and extensive human expertise, with little formal guidance or standardized procedure.
+To address this problem, we propose an iterative model merging framework (Figure 1(c)), leveraging explicit strategies to guide the direction of model evolution toward improved performance.
 We show that even a simple greedy strategy can outperform baseline merging approaches.
-However, in the late stages of both the community experiments and the greedy strategy evolution, achieving further gains in generalization becomes increasingly difficult.
-To explore possible solutions, we introduce ’model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process.
+However, in the later stages of both community-driven model merging experiments and greedy evolution strategies, achieving additional gains in multi-task capability becomes increasingly challenging.
+To explore possible solutions, we introduce ‘model kinship’, a metric inspired by the concept of kinship in evolutionary biology (Sahlins, 2013), to inform and enhance the merging process.
 Model kinship is designed to quantify the degree of similarity or relationship between models throughout the iterative merging process.
-By providing a principled way to measure these relationships, model kinship offers crucial insights for refining merging strategies.
+By offering a principled framework for measuring these relationships, model kinship provides valuable insights that can refine merging strategies for model evolution.
 We conduct a comprehensive analysis of model merging experiments based on model kinship.
-We observe that the community model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau.
-Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of the potential effectiveness of a merge.
+We observe that the model merging process consists of two distinct stages: (1) an improving stage, where models exhibit significant performance gains, and (2) a saturation stage, where improvements diminish and eventually plateau.
+Empirically, we find a strong correlation between model kinship and variations in average task performance, suggesting that model kinship is indicative of potential effectiveness for model merging.
 These findings inspire two main insights: (1) high-kinship merges can lead to performance stagnation, akin to inbreeding; (2) low-kinship merges carry greater risk but may yield larger gains and facilitate escape from local optima.
-Inspired by this, we propose a new continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship.
-[^2] https://github.com/zjunlp/ModelKinship..
+Inspired by this, we propose a novel continual model merging strategy: Top-$k$ Greedy Merging with Model Kinship.
 Specifically, we find that leveraging model kinship as a criterion enables more effective model merging, helping mitigate degradation and avoid local optima during model evolution.
-Model kinship also proves useful as an early stopping criterion, improving the efficiency of the merging process.
-In general, this paper makes four key contributions:
+Furthermore, model kinship also proves useful as an early stopping criterion, improving efficiency of the merging process.
+In general, this paper mainly contains four key contributions:
   1. 1.
-    Continual Model Merging as a Feasible Framework for Model Evolution: We propose continual model merging as a viable framework for evolving large language models.
+    Iterative Model Merging as a Framework for Model Evolution: We propose continual model merging as a viable framework for evolving LLMs.
 Through strategically guided merging across iterations, this approach yields consistent improvements in generalization and task performance.
@@
     Empirical Analysis of Model Evolution: We present a comprehensive empirical analysis of model evolution through iterative merging.
-Our findings highlight the dynamics of multitask performance improvement and stagnation during evolution.
+Our findings highlight the dynamics of multi-task performance improvement and stagnation during evolution.
 In addition, we propose a preliminary explanation of the underlying mechanisms using model kinship.
   4. 4.
-    Practical Model Merging Strategies using Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
+    Practical Model Merging Strategies with Model Kinship: We demonstrate how model kinship guides the model merging process to tackle optimization challenges, and provide practical strategies: Top-$k$ Greedy Merging with Model Kinship, to enhance efficiency and effectiveness of model evolution.
 2 Background
@@
 $\hat{y}=f_{\text{merge}}(x)=\mathcal{F}\left(f_{1}(x;\theta_{1}),f_{2}(x;\theta_{2}),\ldots,f_{n}(x;\theta_{n})\right)$ (1)
-2.2 Continual Model Merging: Benefits and Challenges
-Parameter averaging methods enable the merged model to retain the same architecture and parameter size as the original models, allowing for reuse in future merging processes.
-Leveraging this property, the community has progressively enhanced models through repeated applications of model merging, a process we refer to as “Model Evolution”.
-Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that model evolution can yield highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
-However, current model evolution practices typically involve random merging by multiple contributors, leading to high computational costs and unstable behavior that limit its feasibility for industrial applications.
-3 Strategic Model Evolution
-In this section, we demonstrate how strategic model evolution via continual merging can achieve better generalization across tasks.
-3.1 Method
-We first define an iterative process for model evolution, where models are incrementally merged to produce a more generalized model.
-The direction of model evolution is guided by a principled strategy that governs the model selection.
-At each iteration $t$, a selection strategy $\mathcal{S}$ is applied to the current pool of models $\mathcal{P}_{t}=\{M_{t}^{(1)},M_{t}^{(2)},\ldots,M_{t}^{(k)}\}$, producing a subset of candidate models $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$.
-These selected models are then combined using a merge method $\mathcal{M}$, such as weighted averaging or linear interpolation in the weight space, resulting in the next-generation model $M_{t+1}$.
-This iterative process continues until a predefined end strategy $\mathcal{E}$ is matched.
-Formally, the process is defined as:
-$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{subject to}\quad\neg\mathcal{E}(M_{t+1})$ (2)
+2.2 Model Evolution: Benefits and Challenges
+Parameter averaging methods allow merged models to retain the same architecture and parameter size as their original components, enabling reuse in subsequent merging steps.
+Leveraging this property, the open-source community has progressively enhanced model performance through repeated merging, leading to a “Model Evolution”.
+Empirical evidence from the Open LLM Leaderboard (Beeching et al., 2023) shows that this iterative model merging process can produce highly generalized models, often outperforming those produced through a single merging step (Maxime Labonne, 2024).
+Despite these advances, current community practices largely rely on random merging by multiple contributors, which results in high computational costs and unstable behavior, limiting their practicality for industrial applications.
+3 Iterative Model Merging and Strategy Boost
+In this section, we present controlled experiments to demonstrate that iterative model merging, the process of repeatedly combining models, can help stabilize improvements in multi-task capability.
+Moreover, we demonstrate that incorporating a selection strategy, such as Top-$k$ Greedy, yields more substantial performance improvements.
+3.1 Iterative Model Merging Framework
+We formally define iterative model merging as a modular framework that evolves a population of models through repeated merging.
+It is governed by three components: a selection strategy $\mathcal{S}$, which selects candidate models from the pool $\mathcal{P}_{t}$; a merging operator $\mathcal{M}$, such as SLERP or weighted averaging, to combine the selected models; and an stopping criterion $\mathcal{E}$, which determines when to stop the process.
+At each generation $t$, a subset $\mathcal{S}_{t}\subseteq\mathcal{P}_{t}$ is selected using $\mathcal{S}$, merged via $\mathcal{M}$, and the resulting model $M_{t+1}$ is added back to the pool.
+The process continues until the stopping criterion $\mathcal{E}$ is met:
+$M_{t+1}=\mathcal{M}(\mathcal{S}(\mathcal{P}_{t}))\quad\text{until}\quad\mathcal{E}(M_{t+1})$ (2)
 Method TruthfulQA Winogrande GSM8k Avg.
-Ties 62.76 78.61 0.11 47.16
-Linear 56.37 78.08 68.54 67.66
-Continual Merge (ord1) 47.15 76.24 53.15 58.84
-Continual Merge (ord2) 61.01 79.56 63.76 68.11
-Continual Merge (ord3) 49.80 78.53 55.72 61.35
-Strategic Model Evolution 50.94 80.11 75.13 68.72
-Table 1: Performance Comparison across different merging methods.
+Ties (Density=0.5, Weight=0.3) 62.76 79.56 8.79 50.37
+Dare Ties (Density=0.5, Weight=0.3) 59.36 79.08 65.73 68.05
+Linear (Weight=0.3) 56.37 78.08 68.54 67.66
+Sequential SLERP Merge (ord1) 47.15 76.24 53.15 58.84
+Sequential SLERP Merge (ord2) 61.01 79.56 63.76 68.11
+Sequential SLERP Merge (ord3) 49.80 78.53 55.72 61.35
+Random Merging (k=3) 54.32 78.53 72.81 68.55
+Top k Greedy Merging (k=3) 50.94 80.11 75.13 68.72
+Table 1: Performance Comparison across multi-model merging, sequential model merging with different SLERP orders, and iterative model merging with strategy.
 3.2 Setup
-Baseline Method.
-We adopt two commonly used approaches for merging multiple models as baselines.
-The first is multi-model merging, where all models are merged simultaneously.
-The second is sequential continual merging, where three models are merged in sequence using a pairwise merging strategy.
-For these baselines, we apply TIES (Yadav et al., 2023) and Linear Merge for the multi-model merging case and SLERP (Spherical Linear Interpolation) (Shoemake, 1985) for the pairwise merging in the sequential strategy.
-Top $k$ Greedy Merging.
-Our method applies continual model merging with greedy strategy, the Top-$k$ Greedy Merging approach on $n$ LLMs (as outlined in the black section of Algorithm 1 in Appendix C).
-Each single merging step in the continual merging experiments is performed using SLERP.
+Baseline Methods.
+We consider two types of baselines: (1) Multi-model merging methods that support merging all models at once, such as TIES (Yadav et al., 2023), Dare-TIES (Yu et al., 2024b), and Linear; and (2) Sequential merging, where models are merged pairwise using SLERP (Shoemake, 1985).
+Top-$k$ Greedy Merging.
+Our approach applies iterative model merging using a Top-$k$ greedy selection strategy on $n$ LLMs (see Algorithm 1).
+Each merge step uses SLERP.
+We also include a random merge baseline (Appendix C.2).
 Models and Datasets.
-We select three fine-tuned, open-source LLMs based on the Mistral-7B architecture from HuggingFace: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
+We use three HuggingFace LLMs based on Mistral-7B: mistral-7b-instruct-v0.2, metamath-mistral-7b, and open-chat-3.5-1210.
 Evaluation.
-Evaluation is performed using three task-specific benchmark datasets: Winogrande, GSM8k, and TruthfulQA.
-These benchmarks demonstrate the distinct strengths of the three selected fine-tuned models.
-Further details on the tasks are provided in Appendix A.3.
+Models are evaluated on Winogrande, GSM8k, and TruthfulQA, which highlight their task-specific strengths. Dataset details are in Appendix A.3.
 3.3 Results
-As shown in Table 1, Strategic Model Evolution can yield better generalization when combining multiple tasks.
-However, compared to simple sequential merging, the improvements are limited due to the simplicity of the greedy strategy.
-This observation raises an important question regarding optimal merging strategies:
-Problem: How can we design more principled and effective strategies for model merging to better facilitate model evolution?
+As shown in Table 1, iterative model merging can yield better generalization when combining multiple tasks.
+In particular, both random and Top-$k$ greedy iterative merging outperform one-step baselines, demonstrating the effectiveness of continual merging.
+While single-step merging methods can sometimes achieve strong results (e.g., Linear or Dare Ties), their performance is often unstable and sensitive to hyperparameter.
+Ties, for instance, fails drastically on GSM8k.
+Sequential SLERP merging shows similar limitations, as its performance varies significantly depending on the merge order.
+In contrast, iterative merging strategies are more stable, consistently yielding robust performance across tasks.
+However, these methods offer only marginal improvements. This raises a key question: How can we further enhance performance beyond the current limits of naive or greedy strategies?
 4 Preliminary Analysis of Model Kinship
-To explore efficient strategies for model evolution, we propose a novel metric, model kinship, which captures task-related differences between models and may help predict merging outcomes.
-We then conduct a preliminary analysis of model kinship in community merging experiments on LLMs to evaluate its potential for informing and enhancing model evolution strategies.
+To address this limitation, we investigate whether structural signals beyond raw performance can guide model selection.
+We introduce model kinship, a metric capturing parameter-space similarity, and conduct a preliminary analysis to examine its correlation with merge gains in community-merged LLMs.
+This helps assess the potential of kinship-aware strategies in guiding continual merging and avoiding local optima.
 4.1 Model Kinship
 Drawing inspiration from the parallel between artificial selection and model evolution (as detailed in Appendix G), we hypothesize that a concept analogous to kinship, the genetic relatedness studied in evolutionary biology (Thompson, 1985), can also apply to model merging.
-Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between models.
-This analogy suggests that just as genetic kinship influences breeding outcomes, model kinship may similarly impact the effectiveness of merging strategies in enhancing generalization performance.
+Specifically, we introduce the notion of model kinship, a metric designed to capture and quantify the evolutionary relationships between the merge candidates.
+This analogy suggests that, as genetic kinship affects breeding outcomes, model kinship similarly influences the effectiveness of merging strategies in enhancing multi-task performance.
 We adopt the most intuitive representation, inspired by the cosine similarity analysis introduced in the Task Arithmetic paper (Ilharco et al., 2023).
@@
 For the calculation, $sim(\cdot,\cdot)$ denotes the similarity metric function used.
-Considering two cases merging of 2 models and merging of $n$ models, we formally define model kinship $r$ as:
+Considering two cases merging of 2 models and merging of $n$ models, we define model kinship $r$ as:
 $r=\begin{cases}sim(\delta_{1},\delta_{2}),&\text{(Merge 2)}\\ \frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}sim(\delta_{\text{i}},\delta_{\text{j}}),&\text{(Merge N)}\end{cases}$ (4)
@@
 Each task $T_{i}$ in the set $T$ is associated with a performance measure $P_{i}$ for the LLM.
-For a multitask objective, the Average Task Performance (Avg.) $\bar{P}$ is calculated using the equation:
+For a multi-task objective, the Average Task Performance (Avg.) $\bar{P}$ is calculated by:
 $\bar{P}=\frac{1}{n}\sum_{i=1}^{n}P_{\text{i}}$ (5)
@@
 In this section, we analyze model kinship from two perspectives: (1) its correlation with performance gain across a broad range of open-sourced LLM merges, and (2) its dynamic along specific model evolution paths.
-These analyses aim to clarify the relationship between model kinship and multitask capability improvements, as well as to identify phases of merge effectiveness.
+These analyses aim to clarify the relationship between model kinship and multi-task capability improvements, as well as to identify phases of merge effectiveness.
 4.3.1 Correlation Between Model Kinship and Performance Gain
 We begin by exploring the potential relationship between merge gain and model kinship using three similarity metrics: Pearson Correlation Coefficient (PCC), Cosine Similarity (CS), and Euclidean Distance (ED).
-The models used in this analysis are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
+The models used are based on the Mistral-7B architecture (Jiang et al., 2023a) and collected from HuggingFace, with reference to the Open LLM Leaderboard (see Appendix A).
 As illustrated in Figure 2, the scatter plots derived from all three metrics suggest a moderate correlation between model kinship and merge gain.
@@
 Figure 5: Left (a): The comparison of task performance improvement across merging generations.
-The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multitask capabilities.
-Right (b): The partial model family tree from the controlled experiments.
-The red arrow shows the critical change between experiment 1 and experiment 2 in the evolution path.
+The red curve (greedy strategy) saturates by generation 2, while the blue curve (modified strategy) escapes the local optima at generation 2 and continues improving multi-task capabilities.
+Right (b): The partial model family tree from controlled experiments.
+The red arrow shows critical changes between experiment 1 and 2 in the evolution path.
 Figure 3 displays the average task performance and the merge gains along the two evolution paths. The merging process exhibits two distinct phases:
-  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multitask balance.
+  • Improving Stage. Rapid performance gains and significant merge improvements, driven by active multi-task balance.
   • Saturation Stage. Performance stabilizes, and additional merges result in minimal or no measurable improvement.
 Figure 4 shows how the model kinship and normalized average performance change over the course of the merging process.
-Both metrics display a similar two-stage trend: they increase during the Improving Stage and level off during the Saturation Stage.
-This parallel trajectory suggests a potential link between model kinship and performance gains, indicating that improvements in generalization may coincide with and possibly depend on increases in model kinship.
+Both metrics exhibit a consistent two-phase trend: an Improving Stage, characterized by steady increases, followed by a Saturation Stage, where growth plateaus.
+This parallel progression highlights a potential correlation between model kinship and performance gains, indicating that enhancements in generalization are not only concurrent with but may also be facilitated by increases in model kinship.
 Metric Correlation Correlation
@@
 Table 2: Correlation of Model Kinship based on different correlation function $sim(\cdot,\cdot)$ with Merge Gain, along with their corresponding p-values.
-Additionally, to generalize the findings from the Evolution Paths, we analyze how model kinship evolves across different stages of the merging process, thereby broadening the scope from individual paths to the full evolution picture.
-(detailed in Appendix D.1).
-The results show that the best performing models exhibit high mutual kinship, which may lead to a stagnation stage in the merging process.
-4.4 Discussion
-Considering all results that we observed, this analysis provides two main inspirations for the application of model kinship:
-  • High kinship merges may lead to performance stagnation, similar to biological inbreeding.
-  • Low kinship merges involve risk, but can lead to greater gains, potentially enabling escape from local optima caused by greedy strategy.
 Greedy Strategy + Model Kinship
-Model Avg. Gain Kinship Model Avg. Gain Kinship
-MetaMath 63.72 / / MetaMath 63.72 / /
-Instruct 61.82 / / Instruct 61.82 / /
-Open-chat 66.28 / / Open-chat 66.28 / /
-model-1-1 62.17 -0.6 0.01 model-1-1 62.17 -0.6 0.01
-model-1-2 64.02 -0.03 -0.02 model-1-2 64.02 -0.03 -0.02
-model-1-3 66.84 +1.84 0.05 model-1-3 66.84 +1.84 0.05
-model-2-1 68.72 +2.16 0.93 model-2-1 68.72 +2.16 0.93
-model-2-2 61.47 -3.96 0.57 model-2-2 61.47 -3.96 0.57
-model-2-3 61.32 -3.83 0.58 model-2-3 61.32 -3.83 0.58
-model-3-1 68.59 +1.09 0.95 model-3-2 67.74 +1.09 0.93
-model-3-2 67.74 -0.04 0.93 model-3-3 69.06 +0.74 0.24
-- - - model-3-4 68.61 +1.13 0.32
-model-4-1 68.51 -0.14 0.98 model-4-4 68.75 -0.14 0.54
-model-4-2 68.04 -0.19 0.98 model-4-5 68.39 -0.27 0.66
-model-4-3 68.53 +0.37 0.94 model-4-6 69.03 +0.15 0.52
-- - - model-5-1 69.13 +0.04 0.65
-- - - model-5-2 68.98 +0.07 0.65
-- - - model-5-3 68.63 -0.37 0.98
+[…]
+MetaMath 63.72 / / / MetaMath 63.72 / / /
+Instruct 61.82 / / / Instruct 61.82 / / /
+Open-chat 66.28 / / / Open-chat 66.28 / / /
+model-1-1 62.17 -0.6 -4.11 0.01 model-1-1 62.17 -0.6 -4.11 0.01
+model-1-2 64.02 -0.03 -2.26 -0.02 model-1-2 64.02 -0.03 -2.26 -0.02
+model-1-3 66.84 +1.84 +0.56 0.05 model-1-3 66.84 +1.84 +0.56 0.05
+model-2-1 68.72 +2.16 +1.88 0.93 model-2-1 68.72 +2.16 +1.88 0.93
+model-2-2 61.47 -3.96 -7.25 0.57 model-2-2 61.47 -3.96 -7.25 0.57
+model-2-3 61.32 -3.83 -7.40 0.58 model-2-3 61.32 -3.83 -7.40 0.58
+model-3-1 68.59 +1.09 -0.13 0.95 model-3-2 67.74 +1.09 -0.98 0.93
+model-3-2 67.74 -0.04 -0.98 0.93 model-3-3 69.06 +0.74 +0.34 0.24
+- - - - model-3-4 68.61 +1.13 -0.11 0.32
+model-4-1 68.51 -0.14 -0.08 0.98 model-4-4 68.75 -0.14 -0.31 0.54
+model-4-2 68.04 -0.19 -0.67 0.98 model-4-5 68.39 -0.27 -0.36 0.66
+model-4-3 68.53 +0.37 -0.06 0.94 model-4-6 69.03 +0.15 +0.42 0.52
+- - - - model-5-1 69.13 +0.04 +0.10 0.65
+- - - - model-5-2 68.98 +0.07 -0.15 0.65
+- - - - model-5-3 68.63 -0.37 -0.50 0.98
 Table 3: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach.
@@
 Note: The model kinship experiment was terminated at generation 5, as it has already outperformed the greedy strategy by that point.
+To move beyond individual Evolution Paths, we further investigate how model kinship develops across different stages of the merging process, thereby extending our analysis to the broader evolutionary landscape (refer to Appendix D.1).
+Our findings reveal that the highest-performing models maintain strong mutual kinship.
+However, this close relatedness can also induce a stagnation phase, where the lack of diversity limits further performance improvements.
+4.4 Discussion
+Considering all results that we observed, this analysis provides two main insights for the application of model kinship:
+  • High kinship merges may lead to performance stagnation, similar to biological inbreeding, where excessive similarity limits the ability to adjust.
+  • Low kinship merges involve risk, but can result in greater gains, potentially enabling escape from local optima caused by the greedy strategy.
 5 Using Model Kinship to Improve Model Merging
-Building on the insights from section 4, we explore how model kinship can be leveraged to improve the model merging process.
+Building on the insights from section˜4, we explore how model kinship can be leveraged to improve the model merging process.
 Our main experiment centers on the Mistral-7B model, with detailed results presented in the main text.
@@
 5.1 Main Experiment Setup
-For the main experiments, we follow the same settings as described in section 3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA.
-We adopt the Top-$k$ Greedy Merging strategy as the baseline continual merging strategy.
+For the main experiments, we follow the same settings as described in section˜3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA.
+We adopt the Top-$k$ Greedy Merging strategy as the baseline iterative merging strategy [^2] We provide a Google Colab Notebook..
 Top $k$ Greedy Merging with Model Kinship.
@@
 Figure 5 (a) illustrates the improvements in top average task performance across merging generations.
-Table 3 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
-While both strategies achieve the multitask objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72.
+Table 3 provides model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method.
+While both strategies achieve the multi-task objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72.
 In contrast, strategy utilizing model kinship escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
@@
 To explore how low-kinship merges help escape local optima during saturation, we analyze weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3), with $v_{pre}$ (from Model-1-3 to Model-2-1) as baseline.
-Figure 6 shows that merging with the exploration model ($v_{2}$) induces significant, novel weight changes, while $v_{1}$ shows minimal change due to high similarity between openchat-3.5 and Model-2-1.
+Figure 6 shows that merging with the exploration model ($v_{2}$) produces significant, novel weight changes, while $v_{1}$ shows minimal change due to the high similarity between openchat-3.5 and Model-2-1.
 Early Stopping at High Kinship can Improve Efficiency.
-Iterative merging is resource-intensive.
+Iterative model merging can be resource intensive.
 In our main experiments, a greedy strategy saturated after 2/4 merges with no further gains.
@@
 Since the 2010s, weight averaging has been widely applied in deep neural networks, notably for combining checkpoints to improve training stability and performance.
-(Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Smith and Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022).
+(Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Ilharco et al., 2022; Izmailov et al., 2018; Smith and Gashler, 2017), and parallel training of large language models (LLMs) (Li et al., 2022).
 Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022).
@@
 (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base.
-Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights.
+Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024b), which prevents interference by randomly dropping weights.
 The Model Breadcrumbs (Davari and Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging.
@@
 Recent work like FuseChat (Wan et al., 2024b) integrates it with Knowledge Fusion (Wan et al., 2024a) to enable novel fusion approaches.
+Li et al.
+(2025) further demonstrate that model merging can stabilize training and serve as a low-cost simulator for annealed performance during pretraining, enabling checkpoint reuse and faster validation.
 Recently, there have been some works exploring “model evolution”.
 Tellamekala et al.
-(2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models.
+(2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multi-task models.
 Labonne (2024) develop a tool to automatically merge models on HuggingFace.
@@
 (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations.
+Tang et al.
+(2025) present a continual model merging method that enables training-free projection-based merging of models as they arrive sequentially, improving efficiency and task retention without retraining.
 7 Conclusion
-We propose continual model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance.
+We propose iterative model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance.
 To support this framework, we introduce model kinship, a metric that guides merge candidate selection and explains both performance gains and stagnation during merging.
@@
 Limitations
-However, there are several limitations to consider: a) The experiments in this study are conducted on models with two architectures, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu and Dao, 2023).
-Furthermore, scaling of tasks and candidate models requires further experimentation to understand the computational cost across various scenarios.
-b) The analysis is based on open source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias.
-c) Possible correlation metrics for model kinship have not been fully explored.
-Other metrics may perform better than those discussed in this paper.
-d) The effectiveness of model kinship is demonstrated through empirical evidence.
-However, a theoretical framework (such as the assumptions in Appendix 15) is needed to explain model evolution and model kinship more rigorously.
-e) Model kinship currently guides merging and enhances performance, but does not support sustained evolution.
-Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
+This study has several limitations: a) Our experiments are limited to two model architectures.
+It remains unclear whether our metric and method generalize to other architectures, such as Mamba (Gu and Dao, 2023).
+In addition, the scalability of tasks and candidate models requires further evaluation to assess computational costs in diverse settings.
+b) The analysis relies on open-source data from the Open Leaderboard.
+As this dataset is community-generated, it may be affected by noise or user bias.
+c) We have not fully explored alternative correlation metrics for measuring model kinship.
+Other metrics may yield stronger or more consistent performance.
+d) Our demonstration of model kinship is empirical.
+A more rigorous theoretical framework, such as the assumptions outlined in Appendix 15, is needed to explain model evolution and kinship in greater depth.
+e) While model kinship currently guides merging and improves performance, it does not support sustained evolution.
+Future progress may require incorporating environmental feedback, reward models (Silver et al., 2021), and novel architectures.
+Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No.
+62576307), Ningbo Natural Science Foundation (2024J020), the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00), the Academic Research Fund Tier 1 (FY2025) (Grant T1 251RES2507), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
 Appendix A Details of Experiments in Main Sections
@@
   • Branch Identification: We identified the two longest branches within each tree, representing significant sequences of model merging.
-  • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multitask capabilities and model kinship metrics.
+  • Performance and Kinship Evaluation: These branches were analyzed for changes in merging performance, particularly focusing on shifts in multi-task capabilities and model kinship metrics.
 Table 4 and 5 present detailed information on the sequential merging process.
@@
 Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of models at each stage.
-For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
+For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage (<<0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
 D.2 Details of Model Group Selection
@@
 Appendix E Analysis between Task Relatedness and Model Kinship
-In the formulation of model kinship, we use the placeholder (,$\sim$($\cdot$,$\cdot$)∼ ( ⋅ , ⋅ ) as a similarity metric function to explore options that can effectively capture task-related differences.
+In the formulation of model kinship, we use the placeholder $\sim(\cdot,\cdot)$ as a similarity metric function to explore options that can effectively capture task-related differences.
 One such metric is cosine similarity, derived from the analysis in the task vector, which has been validated as effective for representing differences in single-task models through the cosine similarity of delta parameters (task vectors).
 In addition to cosine similarity, we also investigate the Pearson correlation coefficient and Euclidean distance.
-However, we have not thoroughly evaluated the applicability of these metrics in the context of model evolution, particularly for merged models with multitask capabilities.
+However, we have not thoroughly evaluated the applicability of these metrics in the context of model evolution, particularly for merged models with multi-task capabilities.
 To address this, we examine the relationship between the similarity metrics and task information in subsequent sections.
@@
 Figure 9: An intuitive illustration of how model evolution can fall into local optima due to a performance-prior strategy.
-It shows that Merged Model 2 may be overlooked, despite its potential for better multitask performance.
+It shows that Merged Model 2 may be overlooked, despite its potential for better multi-task performance.
 Our findings using new strategy offer a new perspective on model evolution through multiple merging.
@@
 Figure 10: An intuitive illustration of the optimization process assumption in model evolution, where models progressively converge towards the optimal model.
-Figure 11: An intuitive comparison between selective breeding and continual model merging.
+Figure 11: An intuitive comparison between selective breeding and iterative model merging.
 The left process demonstrates breeding a tall and frutful plant by selecting parents with the desired traits in an biological scenario.